### v1

# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

Abstract
Differential privacy (DP) enables private data analysis but is hard to use in practice. For data controllers who decide what output to release, choosing the amount of noise to add to the output is a non-trivial task because of the difficulty of interpreting the privacy parameter $\epsilon$. For data analysts who submit queries, it is hard to understand the impact of the noise introduced by DP on their tasks.
To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm that chooses $\epsilon$ automatically; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows the controller to control the data flow and achieve trustworthiness while maintaining high performance. We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms. All in all, our work contributes to making DP easier to use by lowering adoption barriers.
1 Introduction
Analyzing private data is essential for various tasks, and differential privacy (DP) is a promising tool to achieve privacy. In a typical DP deployment, data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller. The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals. The data controller wants to protect their identity while releasing useful output to the data analyst. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite bespoke deployments by large organizations such as the US Census Bureau[3], Apple[5], and Google[24], it remains difficult to use DP in practice [23][43]. In this paper, we consider two important challenges controllers and analysts face.
Challenge 1. The controller’s perspective. Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$. Dwork et al.[23] showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
Our contribution. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a query output affects individuals’ privacy. And because the PRI and $\epsilon$ are connected in a principled way, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy requirement. We exploit this opportunity by designing an algorithm that automatically finds an $\epsilon$ and releases query output such that all individuals receive adequate protection based on the controller’s privacy requirement.
Challenge 2. The analyst’s perspective. Analysts submit queries and receive noisy output that deviates from the ground truth output, i.e., the output if DP had not been applied. Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks; this hinders the analysts’ ability to use the DP output in practice.
Our contribution. In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control the data flow. In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform. With the controller’s permission, the platform invokes the evaluation function on both the DP output and the ground truth output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
Putting everything together: Escrow-Based Platform. The utility signaling protocol requires analysts to place more trust in the platform than in a traditional DP deployment. We want to make the platform trustworthy while maintaining high performance, which is at odds with the increased overhead introduced by the algorithm we propose to help controllers. Our solution is to build a DP platform on top of a data escrow, which permits controlling data flow and improving trust with significantly lower overhead compared to other DP platforms.
Summary of related work. Many approaches have been proposed to help controllers choose $\epsilon$ [30] [26] [23] [36] [39] [40] [28] [48] [32] [28], [26] [30] [36] [39] [40], however, none concentrates on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do. In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users[41][25], i.e., the privacy risk when sharing their data, whereas our approach focuses on helping controllers, who can only choose $\epsilon$ if they understand its privacy implications on those end-users.
Exposing error statistics, e.g., error bound, can help analysts interpret the output; such is the approach followed by the US Census [13]. However, such statistics are not sufficient in many cases. Our utility signaling protocol also complements existing approaches [47] to help analysts gain a deeper understanding of the effect of DP on their output.
Lastly, different deployment of DP has received a lot of attention [20] [42] [9] [4] [45] [46] [49] [10] [44] [11]. Our escrow-based implementation hits a new point in the design space between the degree of trust required by participants and the resulting performance.
[…]
The rest of the paper is organized as follows. We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3. Section 4 presents the utility signaling protocol and the escrow-based platform. We finish with evaluation (Section 5), related work (Section 6) and conclusions in Section 7.
2 Privacy Risk Indicator (PRI)
In this section, we present the new privacy risk indicator. We first provide background on differential privacy.
Differential privacy (DP) [19][21] is a mathematical definition of privacy that bounds the effect a single individual can have on the output of a query. Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O).$
The parameter $\epsilon$ is a tunable parameter called privacy loss budget. A typical way of achieving differential privacy is by adding noise to the output. The amount of noise is controlled by $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output. Therefore, $\epsilon$ is the key parameter for navigating the privacy-accuracy trade-off in differential privacy.
We present two important properties of differential privacy:
Post-processing. [22] DP is immune to post-processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees. Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism. For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
Sequential composition. [22] Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy. Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
2.1 Overview
A good $\epsilon$ provides adequate privacy protection to all individuals while making the output still useful to analysts. However, it is inherently difficult to reason about the privacy implications of choosing $\epsilon$ for an specific individual in the dataset. Because $\epsilon$ bounds the privacy loss for any individual in expectation, but its concrete effect on a specific individual’s privacy is unknown. This is the source of difficulty for controllers whose ultimate goal is to protect individuals in the dataset.
We observe that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to others. For instance, a controller can set an $\epsilon$, compute the corresponding query output, and observe the resulting PRI for each individual. If the PRIs are equal, then all individuals will be equally protected. If they look different, that is an indication that certain individuals are at a higher risk of being identified and a suggestion for the controller to try a different $\epsilon$. The direct connection between the new PRI and differential privacy that we explain next is key to both providing a principled understanding of the new indicator and designing strategies to find $\epsilon$, including an algorithm that finds $\epsilon$ automatically, which we will present in Section 3.
2.2 Privacy Risk Indicator: Formal Definition
We represent a dataset as a set of $n$ records $x=(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, drawn from a data universe $\mathcal{X}$. WLOG, we assume each individual (labeled $1$ to $n$) contributes exactly one record to $x$.
We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. We consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
Definition 1 (Privacy Risk Indicator).
$PRI_{i}=\lVert r-q(x_{-i})\rVert_{1},$
where $x_{-i}$ is the neighboring dataset of $x$ with individual $i$’s record $x_{i}$ removed. The difference is in the L1 norm.
Computing PRI is $\epsilon$-DP. Since $r$ is the output of an $\epsilon$-differentially private mechanism, computing PRI also satisfies $\epsilon$-differential privacy by DP’s post-processing property.
Connection between PRI and privacy loss. The PRI has a direct connection to the concept of privacy loss as defined in differential privacy. We formally show that when the Laplace Mechanism is used to compute the query output $r$, the connection is simple: larger PRI corresponds to larger privacy loss and vice versa.
First, we introduce the definition of privacy loss in DP[22].
Definition 2 (Privacy Loss).
Given two neighboring dataset $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
$\mathcal{L}_{\mathcal{M}(x,q,\epsilon)\|\mathcal{M}(x^{\prime},q,\epsilon)}^{(r)}=\ln\left(\frac{\Pr\left[\mathcal{M}(x,q,\epsilon)=r\right]}{\Pr\left[\mathcal{M}(x^{\prime},q,\epsilon)=r)\right]}\right).$
Next, we introduce the definition of Laplace Mechanism[22], a classic $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
Definition 3 (Laplace Mechanism).
Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon>0$, the Laplace Mechanism is defined as:
$\mathcal{M}_{L}(x,q,\epsilon)=q(x)+(Z_{1},\ldots,Z_{k}),$
where $Z_{i}\sim Lap(\frac{GS_{q}}{\epsilon})$ are i.i.d random variables drawn from the Laplace distribution  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
Using the Laplace Mechanism, the relative privacy loss of a specific individual $i$ can be expressed as:
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}$ $=\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/GS_{q})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/GS_{q})}\right)$
$=\frac{\epsilon(\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1})}{GS_{q}}.$
Since $\lVert r-q(x_{-i})\rVert_{1}$ is the definition of $PRI_{i}$, there exists a clear connection between $PRI_{i}$ and individual $i$’s privacy loss: larger PRI corresponds to larger privacy loss. Similarly to PRI, the privacy loss for a specific individual is data and query-dependent; essentially, it is the privacy loss individual $i$ will suffer if they participate in the analysis and the corresponding query output $r$ is released, compared to if they do not participate. Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss will always be bounded by $\epsilon$.
We have shown that the PRI indicates the relative privacy loss of each individual if a given output is released using DP. A data controller can leverage the PRI for many purposes; one natural goal is to exploit the relationship between PRI and DP to choose an $\epsilon$ and the corresponding output to release. In the next section, we explain different approaches the controller can use to achieve this goal and propose an algorithm.
3 Deriving Epsilon based on PRI
In this section, we describe how to leverage the PRI we introduced to address Challenge 1. We explain the two general approaches of using the PRI to find $\epsilon$ in Section 3.1. We describe the key observations we made to determine whether the PRIs offer adequate protection and the detailed algorithm for choosing $\epsilon$ and releasing query output in Section 3.2. In Section 3.3, we consider multiple queries and discuss how to bound the total privacy loss. Finally, we discuss the algorithm complexities and extensions in Section 3.4.
3.1 Equal PRI means Equal Protection
The goal of DP is to protect each individual equally in expectation, such that no individual suffers additional privacy loss than others by participating in the analysis; the privacy loss is bounded by $\epsilon$. The PRI indicates the effect of $\epsilon$ on each individual, thus becoming an actionable indicator that aids in selecting a suitable value. We describe two approaches to help controllers choose $\epsilon$ that leads to exactly equal or statistically similar PRIs. We show that both are impractical, which motivates our new algorithm (Section 3.2).
Figure 1: Example of using DP to equalize the PRIs. The goal is to find the largest $\epsilon$ that can make the PRIs equal.
Can we make the PRIs equal? Figure 1 illustrates this using an example dataset with four records. If the output was released directly without adding noise (thus no DP), some individuals would be more at risk than others: those with higher PRI, as shown in the leftmost table. By decreasing $\epsilon$, more noise is added to the output, so the PRIs become more similar. Then, to protect all individuals equally, the controller should choose the $\epsilon$ that makes the PRIs equal. Continuing with the example of Figure 1, setting $\epsilon$ to $0.5$ leads to equal PRIs among the four individuals while $\epsilon=1$ does not. When $\epsilon=0.2$, since it produces larger noise, the PRIs will remain equal, but it also leads to less accurate output. Therefore, $\epsilon=0.5$ is the best choice among the three values as it achieves the best privacy-accuracy trade-off.
Although it helps as an illustration, making PRIs exactly equal across individuals is generally infeasible unless an excessive amount of noise is added to the query output, which is undesirable because this means the output will not be useful for the analyst. It is also possible that a non-zero $\epsilon$ does not exist for which all PRIs are equal, and the only solution is to add an infinite amount of noise and make the PRIs all infinity.
Making PRIs statistically similar? One strategy is to consider the PRIs similar enough if they are statistically similar. Concretely, we can take two biased samples of PRIs: the set of the highest PRIs and the lowest PRIs. Then, we test the null hypothesis $H0$: the two samples come from the same distribution, using the statistical hypothesis framework. We consider the PRIs to be similar enough if we fail to reject $H0$.
Since we choose to compare two biased samples, any hypothesis tests that are sensitive to differences in locations (e.g., Mann-Whitney[37], Kolmogorov-Smirnov[29] or Mood’s Median Test) do not work well in practice since the ranks or the medians of the two samples will likely be very different, thus the tests will always tend to reject the null hypothesis. On the other hand, if we change the strategy to choosing random samples of PRIs and compare them, it becomes too easy to conclude the samples are similar. Therefore, an alternative strategy is needed to determine whether the PRIs provide adequate protection.
3.2 The Find-$\epsilon$-from-PRI Algorithm
Two key observations we make lead to our strategy for determining whether the PRIs provide adequate protection.
O1: Compute ratio between two PRIs. The first observation is we should compute the ratio between the PRIs of two individuals to determine whether they receive similar protection, since the comparison depends on the magnitude of the PRI. Consider two cases: two individuals with PRI of 2 and 1, and another two individuals with PRI of 10 and 9. In both cases, the difference between the two PRIs is 1, but the latter is preferable for equal protection since the ratio is close to 1.
Figure 2: Left tail with similar larger PRIs (green line) is better than right tail with similar smaller PRIs (red line)
O2: Left tail is better than right tail. The second observation is that if we cannot make the PRIs similar to each other as adding too much noise will make the output useless, it is better only to have left tail while making the larger PRIs similar than have right tail; because right tail means some individuals will have much higher risk than the rest, as shown in Figure 2. Even though the green line shows more individuals have relatively larger PRIs, their PRIs are similar enough so that they still receive similar protection; that is, they become indistinguishable with similar privacy loss.
These two key observations lead to a simple strategy for determining whether the PRIs provide adequate privacy protection. Based on O1, we can compute the ratio of the maximum PRI over the minimum PRI. Intuitively, as the ratio approaches 1, the PRIs become more similar and closer to the blue line shown in Figure 2. However, in many cases, this requires adding too much noise. Therefore, O2 allows the controller to trade privacy for more accuracy: instead of choosing the minimum PRI, the controller can choose the PRI at a specific percentile $p$. We fix the maximum PRI as the numerator to ensure the riskiest individuals will always be protected. The smaller $p$ is, the larger the ratio; therefore, more noise must be added (smaller $\epsilon$) to bring the ratio closer to 1. Although $p$ is an additional parameter the controller needs to set, it is more interpretable than $\epsilon$, since the controller only needs to decide at which point all the larger PRIs should be similar. For example, if the controller chooses the median (the 50th percentile), a ratio close to 1 indicates no individual is at considerably higher risk than the median individual.
Algorithm. Since PRI and $\epsilon$ are connected in a principled way, we can find a suitable $\epsilon$ that is consistent with the controller’s privacy requirement by computing the PRIs for different $\epsilon$, and choose the $\epsilon$ that is associated with PRIs such that the ratio between the maximum PRI and the PRI at the $p$th percentile is close to 1. The main intuition is to test different values of $\epsilon$ in descending order, and find the largest suitable $\epsilon$ that satisfies the controller’s privacy requirement while preserving as much accuracy as possible (note Figure 1).
Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; DP mechanism $\mathcal{M}$; consumed privacy budget $\epsilon_{c}$; percentile $p$; threshold $\tau$
Output : Chosen privacy parameter $\epsilon$, query output $r$
1 /* Initialize set of $\epsilon$ values to test; minimum value is $\epsilon_{c}$/
2 $\mathcal{E}\leftarrow$ initialize-eps-to-test($\epsilon_{c}$)
3 $x^{p}\leftarrow\textsc{project-query-attrs}(x,q)$
4 /* Compute $q(x_{-i})$ for all $i$ using memoization */
5 $C\leftarrow map()$, $L\leftarrow list()$
6 for $x^{p}_{i}\in x^{p}$ do
7       if $x^{p}_{i}\notin C$ then
8             $C[x^{p}_{i}]\leftarrow q(x^{p}_{-i})$
9      $L.add(C[x^{p}_{i}])$
10for $\epsilon\in\mathcal{E}$ do
11       /* Compute query output using DP with current $\epsilon$/
12       $r\leftarrow\mathcal{M}(x,q,\epsilon)$
13       /* Compute PRI for each individual */
14       $PRI\leftarrow list()$
15       for $l\in L$ do
16             $PRI_{i}\leftarrow\lVert r-l)\rVert_{1}$
17             $PRI.add(PRI_{i})$
18      $PRI_{max},PRI_{p}\leftarrow\max(PRI)$, get-percentile$(PRI,p)$
19       /* A suitable $\epsilon$ is found when the ratio is close to 1 */
20       if $PRI_{max}/PRI_{p}<1+\tau$ then
21             $\epsilon_{c}\leftarrow\epsilon_{c}+\epsilon$
22             return $r$
23
Algorithm 1 Find-$\epsilon$-from-PRI
Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ can be infinite, to ensure the algorithm runs in bounded time, we set a range (e.g., from $\epsilon_{c}$ to 10) and include representative $\epsilon$ value in fixed intervals (e.g., 0.1). The lower bound $\epsilon_{c}$ starts at 0 and updates for each answered query (we explain how to handle multiple queries in Section 3.3). The algorithm then projects relevant query attributes from the dataset (line 1) and computes the query output for each neighboring dataset (line 1 - 1). Next, the algorithm iterates over the list of $\epsilon$ from largest to smallest (line 1) and stops when it finds the first suitable value. For each $\epsilon$ it tests, the algorithm computes the current query output $r$ by applying the given DP mechanism $\mathcal{M}$ with parameter $\epsilon$ (line 1), and use it to compute the corresponding PRI for each individual (line 1 - 1). Finally, the algorithm computes the ratio of the maximum PRI over the PRI at the given $p$th percentile (line 1). If the ratio is close to 1 within the threshold $\tau$ (we use 0.01 in our implementation), then a suitable $\epsilon$ has been found. The algorithm updates the lower bound $\epsilon_{c}$ and returns the DP output $r$ (line 1 - 1).
Algorithm data flow. The algorithm is used by controllers to find $\epsilon$ from the PRIs and release DP output. The controller runs the algorithm on their input dataset and query. The algorithm first computes the PRIs, finds a suitable $\epsilon$ based on the controller’s privacy requirement, and returns the DP output computed using this $\epsilon$. Nothing is released by default to analysts. The PRIs are a by-product of which the sole purpose is to help the controller find $\epsilon$, and they are never released to the data analysts (we will discuss how such promise is enforced in Section 4.3. The $\epsilon$ value may need to be released on certain occasions; since finding $\epsilon$ is a post-processing step to computing PRI, which is DP, its release is also DP.
DP guarantee. The algorithm satisfies $\epsilon$-differential privacy. The algorithm only returns the query output $r$ when a suitable $\epsilon$ can be found. The output $r$ is computed using a standard $\epsilon$-differentially private mechanism (i.e., Laplace Mechanism); thus, the algorithm shares the same DP guarantee. This is no different than a standard DP deployment, only that $\epsilon$ is chosen automatically instead of manually by the controller.
3.3 Bounding Privacy Loss Across Queries
The challenge of running multiple queries. In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery [22]. In today’s deployment of DP, the data controller needs to set a total privacy loss budget a priori on the dataset and stop accepting queries when the cumulative $\epsilon$ exceeds the budget. While the total budget provides a rigorous bound of privacy loss in the context of DP, it abstracts away from the actual privacy outcomes, and setting it without context is extremely challenging [23]. To further exacerbate the problem, the data analyst needs to decide the queries they want to run a priori since they need to decide how to utilize the total budget for their queries. This poses a major obstacle for analysts who do not have a pre-determined set of queries.
The solution. Instead of quantifying the maximum acceptable privacy loss, which is hard, we keep track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$. For an incoming query, when using Algorithm 1 to find $\epsilon$, we truncate the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). We reject answering this query if we fail to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$, meaning the PRIs cannot offer adequate privacy protection for any possible $\epsilon$.
Our solution creates a dynamic bound of total privacy loss. As the cumulative privacy loss, $\epsilon_{c}$, keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since less noise can be added to make the PRIs meet the controller’s requirement. At each point that we fail to find a suitable $\epsilon$ for a query, the total privacy loss is bounded. Using DP’s sequential composition property, releasing the output of those answered queries satisfies $\epsilon_{c}$-differential privacy. The bound is refreshed if a suitable $\epsilon$ is found for a different query. This resulting dynamic bound helps controllers bound the total privacy loss without the need to set the total budget a priori. Since the algorithm automatically finds $\epsilon$ (or rejects queries) based on this dynamic bound, it also removes the need for the analyst to decide how to distribute their privacy budget across different queries a priori.
3.4 Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other. We implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 5.3.1 that the algorithm is both efficient and scalable.
Expressing privacy preferences. A modification of our algorithm permits the controller to specify protection preferences over certain groups of individuals whose privacy should be protected similarly. For example, they can specify individuals whose ages are above 60 and those younger need to receive the same privacy protection. This feature can easily be incorporated into the algorithm by letting the controller change the ratio to compute. In the example, the controller can choose to compare the median PRI of the two age groups and choose the $\epsilon$ that makes the ratio of the two medians close to 1.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the data controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp[28] and ViP[40] have explored different ways to visualize privacy-accuracy trade-off in differential privacy. The PRIs can complement the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
4 Helping Analysts Interpret Utility
In this section, we first propose a simple utility signaling protocol to help analysts understand the impact of DP on their downstream applications to address Challenge 2 (Section 4.1). Then, we assemble a set of requirements for a DP platform, which we derive from the solutions proposed (Section 4.2). In Section 4.3, we introduce a new implementation of our new algorithm and protocol on top of a data escrow that meets those requirements.
4.1 The utility signaling protocol
Today’s DP deployments. In the central DP model, we differentiate among individuals who own the original data, controller that manages individuals’ data on their behalf, platform the infrastructure where computation takes place, and analyst the actor demanding computation over individuals’ data. The analyst submit queries to a platform operated by a controller. The platform computes query output using DP and releases them to the analyst with the controller’s approval.
We consider the case where the analyst has a downstream task with an associated evaluation function, g, that returns the utility of the downstream task. The function takes the set of query output, S, that the analyst received as input. Following the notation used in Section 2.2), $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n}\}$, $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$. For example, an analyst queries census data at the county level (i.e., the US Census Bureau is the controller and citizens are the individuals). The set of output $S$ are aggregate statistics on population count, education level, housing prices, etc. The analyst’s task is to build a machine learning model to predict the median income of each county based on those statistics, and the evaluation function returns the prediction accuracy, which is g(S). Since the queries are computed using DP, releasing $g(S)$ is also DP by post-processing.
The problem. The controller uses DP to release noisy output to protect the privacy of individuals, but the analyst does not know how much such noise affects their downstream task and whether the output is useful. In the example, an analyst who obtains a prediction accuracy of $88\%$ does not know how much the accuracy would have been without noise; thus, they cannot decide whether it is worth investing more effort in improving the accuracy of a model. Clearly, the platform cannot tell the analyst how much noise was added because that would jeopardize the privacy guarantee.
A proposed solution: a utility signaling protocol. We propose a protocol to help analysts interpret the implications of the added noise on the quality of their downstream task by letting them compare to the quality they would have obtained over the ground truth output. Let $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n}\}$ be the set of ground truth output (i.e., without applying DP). The protocol allows the analyst to obtain both g(S) and g(S’), thus signaling the utility loss due to DP.
Sending the ground truth output to the analyst so they run their downstream task is not feasible as it breaks the premise of protecting the data in the first place. Instead, the analyst sends their downstream task to the platform. The platform invokes the evaluation function $g$ over both the set of noisy output $S$ and the set of ground truth output $S^{\prime}$. Then, if the controller approves, both $g(S)$ and $g(S^{\prime})$ are released to the analyst to signal the loss in utility due to DP. In the example, the platform trains a model over both the DP census statistics and the ground truth statistics, then applies the evaluation function the analyst shared, obtaining $88\%$ and $90\%$ accuracy, respectively. Upon receipt of both numbers, the analyst understands the utility effect of DP and can decide whether it is worth investing more effort in improving the model.
Assessing Privacy Implications of Releasing $g(S^{\prime})$: Any information release may leak information about the individuals’ data, e.g., even the accuracy of a model (consider the trivial case where $S^{\prime}$ contains only one value, the analyst can simply use the evaluation function to test different values until they obtain an accuracy of 100%). Since DP is not being applied when releasing $g(S^{\prime})$, it is ultimately at the discretion of the controller to decide whether or not to release $g(S^{\prime})$. In the example, the controller may be comfortable releasing the accuracy of the machine learning model trained on the ground truth census statistics (in addition to the model accuracy on the DP statistics), but not releasing the model itself. Our utility signaling protocol does not attempt to address the possibility of various attacks the analyst may perform to infer individual data from $g(S^{\prime})$. Instead, the escrow-based platform we describe next allows the controller to examine the analyst’s downstream task and associated evaluation function and ensures $g(S^{\prime})$ will not be released without the controller’s permission. For many applications, the benefit of allowing the analyst to observe $g(S^{\prime})$ can be tremendous.
4.2 Platform Requirements
We propose a new platform to support the Find-$\epsilon$-from-PRI algorithm that helps controllers (addressing Challenge 1) and the utility signaling protocol that helps analysts (addressing Challenge 2). Such a platform must meet the following requirements:
  • R1. Control Data Flow: Analysts should never see the raw data or PRIs. And they should not submit queries and downstream tasks over the raw data directly without the controller’s approval (e.g., the controller may approve releasing the model accuracy to signal utility loss).
  • R2. Trust: The new protocol requires more trust on the platform. Trust is a broad concept, and its meaning varies depending on context[7]. In the context of DP, trust means the platform adheres to prespecified behavior and the DP mechanism upholds its privacy guarantee[7][43]. Individuals must trust the controller will protect their data in a responsible manner (i.e., using DP). Controllers must trust the platform keeps the data secure and runs computations on the data correctly. With the new utility signaling protocol, analysts must also trust the platform runs their downstream task correctly. We want to reduce the trust those entities have on the platform.
  • R3. High Performance: The platform must limit the overhead, which is at odds with the pressures of the new Find-$\epsilon$-from-PRI algorithm that helps controllers at the cost of higher computational overhead.
We highlight a tension between R2 and R3. Improving trust often requires employing techniques (e.g., cryptographic primitives) that increase the computational overhead, potentially leading to impractical solutions. We now explain the design of a platform that addresses these three requirements.
4.3 Escrow-Based DP Platform
We propose a new platform, based on an escrow architecture similar to the proposed in  [50], that lets controllers maintain fine-grained control of data flow (meeting R1), reduces the trust on the controller and platform by both individuals and analysts (meeting R2) all while maintaining high performance to meet R3.
4.3.1 Controlling Data Flow to Implement the utility signaling protocol
A data escrow supports delegated computation, which means the escrow runs computation on behalf of the computation owner. Implementing a platform with an escrow lets analysts delegate their downstream tasks to the platform, thus obtaining the two utilities $g(S)$ and $g(S^{\prime})$. We present the end-to-end protocol in Figure 3, which consists of seven steps.
Figure 3: End-to-end utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
1. Registering data. The controller registers the dataset containing individual records with the escrow.
2. Registering application. The controller registers an application that uses DP to run queries on the registered dataset, which includes the Find-$\epsilon$-from-PRI algorithm. Any derived data product, such as the PRI, will not be released without the controller’s permission. The PRI can also be computed separately as an independent function inside the application if the controller wants to use it for other purposes.
3. Creating policy. The data controller creates a policy to let the analyst run queries on the registered dataset using the registered application; a policy in the escrow indicates who can run what computation on what data.
4. Issuing Queries. The analyst issues queries to the application. The application answers each query and stops when a suitable $\epsilon$ can no longer be found. After checking that the query computation only ran on data accessible to the analyst, the escrow releases the set of DP output $S$ to the analyst because the controller wrote the corresponding policy. The set of ground truth output $S^{\prime}$ is registered inside the escrow since it is a derived data product of the computation.
Up to this point, the protocol provides analysts with the same information they can access in a typical central DP deployment. Now, we extend the protocol to implement the utility signaling component to help analysts understand the impact of DP on their downstream tasks.
5. Registering downstream task. The analyst registers their downstream task and an associated evaluation function $g$ as an application in the escrow. The evaluation function is the only function within the application that is available for the escrow to invoke and anyone to create policy on; if exists, the rest of the functions should be invoked from the evaluation function to perform other necessary computations. The escrow ensures that only the evaluation function’s output (and nothing else) will be released to the analyst.
6. Creating policy for downstream task. After inspecting the application submitted by the analyst and determining it is acceptable to release $g(S^{\prime})$, the controller creates a policy to run the evaluation function $g$ on the ground truth output $S^{\prime}$.
7. Evaluating downstream task. With the controller’s permission, the analyst obtains both $g(S)$ and $g(S^{\prime})$. By inspecting the difference between the two utilities, the analyst learns the impact of DP on their task, thus helping them decide whether the DP output is sufficiently good for their task.
By exploiting the delegated computation features of the data escrow, the platform controls data flow, which is necessary to implement the utility signaling protocol.
4.3.2 From Trusted to Trustworthy, while Fast
A trusted platform adheres to prespecified behaviors[7], in the context of DP this means an adversary only has access to the output of a DP mechanism, but not any of its internal state and any raw or derived data[22]. This is infeasible to achieve in practice. A central platform, as in the central DP model, can claim to be trusted, but such claim does not preclude it from being compromised and intentionally leaking data in the clear, thus breaking the DP guarantee[43]. A trustworthy platform, on the other hand, implements different techniques to enforce the promise that not only will it never behave unexpectedly, it will not pose harm (to the users) in any way[27]. An escrow-based platform reduces the impractical assumptions of a trusted platform and supports trustworthy computation by ensuring:
  • No data or derived data product (e.g., query output) ever leaves the platform without explicit permission from the controller, and only explicitly allowed computation runs on existing data; this is the basis of the control of data flow.
  • Data is encrypted end-to-end, even during computation. End-to-end encryption can be achieved in many ways; for instance, DP platforms leveraging secure multiparty computation[42][9][10][49][45], homomorphic encryption[4][46], Oblivious RAM[12] and trusted execution environment[44]. The prototype we use for our evaluation uses AMD Secure Encrypted Virtualization(SEV)[33].
  • Tasks that are sent to the escrow are the tasks that run. The escrow is registered with an escrow application; the platform ensures that the correct version of the application is running using techniques such as remote attestation[15] or formal verification[38][35]. Our prototype uses hardware-rooted remote attestation enabled by AMD SEV.
Figure 4: Left: central platform. Right: escrow-based platform
By computationally enforcing the promises above, the escrow reduces the trust that both individuals and analysts must place in the platform while maintaining the same accuracy guarantee as the central DP model. Furthermore, the escrow application can be written so that controllers do not gain access to the raw data but only the PRIs (i.e., by writing a policy that only reveals the PRIs to the controller), thus illustrating multiple escrow-based deployment scenarios. Consider Figure 4, which shows how individuals, controller, and analyst interact with the central platform in CDP and the escrow-based platform. In CDP, the controller manages individuals’ data using the central platform and sets $\epsilon$ manually; the individuals need to trust both the controller to protect their privacy (i.e., by choosing a good $\epsilon$) and the platform to operate correctly without leaking their data. Whereas using the escrow-based platform, individuals send data directly to the platform, and the controller only has access to PRIs computed by Algorithm 1. The escrow invokes Algorithm 1 to find $\epsilon$ on the controller’s behalf and releases the output to the analyst, who delegates their downstream task to the escrow to obtain utility.
5 Evaluation
Our evaluation shows our solutions address both Challenge 1 to help controllers, and Challenge 2 to help analysts. We answer the following research questions for our evaluation:
  • RQ1: Does presenting the privacy risk indicator help data controllers choose $\epsilon$?
[…]
  • Microbenchmarks: We conduct experiments to show the
    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study how sensitive is our algorithm to different values of parameter $p$.
5.1 RQ1: Using PRI To Choose Epsilon
To determine whether the PRI help the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$. Our goal is to show that by presenting the PRI to the participants, they will be more consistent on the $\epsilon$ they choose compared to if they don’t have the PRI. To this end, we conducted a between-subject user study where the participants are divided into control and treatment groups. To prevent learning effects, only the treatment group sees the PRI. We chose to conduct the study in the form of an online survey, because our tasks do not require in-person guidance and it is more convenient for participants to complete the study. In addition, it allows us to collect data from a more representative sample.
5.1.1 Recruitment and data collection
We recruited 100 participants using Prolific[1]. To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP Address, location data, and contact information are not being recorded.
Prescreening criteria. When recruiting on Prolific, we prescreen participants based on three criteria: i) they are fluent in English, to avoid confusion due to misunderstanding language; ii) they have completed an undergraduate degree or above; iii) they study Computer Science, Computing (IT) or Mathematics; this is to ensure that the participants have a sufficient quantitative background to complete the study.
Tutorial and Screening Quiz. Since the main goal of the study is to ask the participant to choose $\epsilon$ when using DP to release query output, the participant needs to have basic knowledge about DP and understand the implications of the $\epsilon$ they chose. To ensure each participant knows enough about DP to complete the study, we ask each participant to read a short tutorial about DP at the beginning. This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level. Mainly, it explains DP is generally achieved by adding random noise to the query output, and $\epsilon$ is inversely proportional to the amount of noise added.
After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read. Only participants who pass the quiz proceed to conduct the study. Out of the 100 participants we recruited, 56 of them passed the quiz and their responses are used for the study analysis. This quiz ensures that participants have understood the principles of DP and thus ensures the validity of the user study.
5.1.2 Study procedure
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group and 27 participants are in the treatment group. The survey presents a sample from the UCI Adult dataset[2] and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order and asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 and each interval being 0.5. The participant’s task is to choose an $\epsilon$ value that offers equal privacy protection to all people in the dataset while preserving as much accuracy as possible.
For the control group, the participant is presented with a table that shows the range of possible output for each $\epsilon$, indicating output accuracy, along with the ground truth (non-DP) output. The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
For the treatment group, the participant is presented with the same table, but with an additional column that includes a plot showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI). Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. We explain beforehand what the PRI is at a high level, and that a smaller range represents more similar PRIs among people in the dataset. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
5.1.3 Results
Figure 5: $\epsilon$ chosen by each participant
We compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups. Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control and treatment group for each query. For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group. This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy. This shows that most participants could connect the PRIs and their task goal, which is to offer equal privacy protection while ensuring high accuracy, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal. On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to pick.
In addition, we ran the Mann–Whitney test [37] on the $\epsilon$ values in the control and treatment group, with the alternative hypothesis being the two samples come from different distributions. For both queries, we obtained $p<0.005$. Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
In the follow-up question for the treatment group, all 27 participants answered yes when asked whether they used the PRIs when choosing $\epsilon$, confirming our hypothesis that presenting the PRIs helped participants choose the $\epsilon$ that satisfies the task goal.
Figure 6: $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
Does knowing DP matter? Another interesting question is whether the participant’s a priori familiarity with DP affects how they choose $\epsilon$. At the start of the survey before showing the tutorial, we asked each participant whether they know about DP. 27 participants (out of a total of 56) answered they don’t know anything about DP and 29 participants answered they know about DP.
Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity. We find similar patterns. Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance. This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they know about DP previously.
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$, instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs - by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
Figure 7: $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference. Participants who saw histograms did not perform better than participants who saw range plots. Especially for Q2, 8 participants chose larger $\epsilon$ (3.5, 4, and 4.5). Although we cannot be certain since the responses are anonymous, one possible explanation might be that for the large $\epsilon$, the majority of PRI are already small and similar to each other. Only a small portion of PRI have large values. The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate. However, if the participants only saw the range of PRI, they would be more inclined to choose smaller $\epsilon$ with a smaller range of PRI. Another possible explanation could be that some participants had difficulties interpreting histograms, whereas the range plots are generally much easier to understand.
The discrepancy of results by tweaking how we present the PRIs demonstrates how we convey information to the users affects their ability and reasoning for choosing $\epsilon$. Exploring other alternative solutions for presenting and helping users interpret PRI remains an interesting problem.
5.1.4 Threats to validity
External validity. We recruited participants from diverse backgrounds across the world via Prolific, with few prescreen criteria such that the participants can complete the study without difficulty.
Internal validity. To ensure all participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz. We only allowed the participant to take the quiz once to minimize the chance of participants abusing it if they got multiple chances. To avoid learning effects, we used a between-subject study to isolate the condition we intended to test. We also randomized the two questions for choosing $\epsilon$.
5.2 RQ2: DP On Escrow-based Platform
[…]
5.2.1 Baselines for comparison
We evaluate and compare the performance of the four following baselines:
  • Data Escrow: We implemented a private query escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output. Optionally, if the parameter $\epsilon$ is not given, it will use Algorithm 1 to find $\epsilon$ automatically. We simulated the protocol where the data controller registers their dataset to the escrow, creates a policy that allows the data analyst to issue queries to their data using the private query application, and the analyst issues their query and receives the query output.
[…]
  • Standalone without DP: This baseline simply runs the given query without DP. The dataset is in plaintext.
  • Standalone with DP: This baseline represents the central DP model. It runs the given query and applies the Laplace Mechanism to the query output. The dataset is in plaintext.
[…]
5.2.2 Setup
We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine. We used AMD’s Secure Encrypted Virtualization (SEV) when running the data escrow. We used Ubuntu 20.04 and Python 3.8 for all baselines.
5.2.3 Dataset
[…]
5.2.4 Queries
We run the 3 queries independently for all baselines:
  • Q1: Number of people between age 20 and 30 (simple predicate with range)
  • Q2: Number of people by each race (aggregate & group-by)
  • Q3: Number of people who are female and have income greater than 50K (two simple predicates)
5.2.5 Results
[…]
Figure 8: Average query time over 10 runs
[…]
5.3 Microbenchmarks
We evaluate the runtime and scalability of the Find-$\epsilon$-from-PRI algorithm in Section 5.3.1 and study how different values of the parameter $p$ affect the $\epsilon$ chosen in Section 5.3.2.
5.3.1 Is Find-$\epsilon$-from-PRI practical?
Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical. We ran the algorithm as a registered application in the escrow.
Datasets. We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset. We sampled records for size datasets of size 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they are not the same copy of each other. We included all attributes and all records in the dataset.
Queries. We ran the algorithm to find $\epsilon$ for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.
  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num=13 AND age==25
  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==’Asian-Pac-Islander’ AND age>=30 AND age<=40 GROUP BY marital_status
  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=’United-States’ AND sex==’Female’
  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass==’Federal-gov’ OR workclass==’Local-gov’ or workclass==’State-gov’
  • Q5: SELECT SUM(fnlwgt) FROM adult WHERE capital_gain>0 AND income==’<=50K’ AND occupation==’Sales’
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,0.099,\cdots,0.001]$ (37 values in total). We set $\epsilon_{c}=0$, $p=50$ and $\tau=0.01$. We used six parallel processes for computing the PRIs.
Figure 9: Average runtime to find $\epsilon$ over 10 runs
Results. We first note that all queries on all 10 runs were able to find a suitable $\epsilon$ on every dataset we tested. Figure 9 shows a line plot where the x-axis is the data size and the y-axis is the runtime, both in log scale. Each line represents a query, with the error band showing the standard deviation.
As shown in the figure, the runtime increases linearly with the data size. Most queries took less than a few minutes to finish, even for 1 million records, except Q5, which took more than an hour. The runtime for Q5 is slower since it involves the attribute fnlwgt, which is a numerical attribute with a large domain size and many unique values, thus the query involves significantly more unique records than the other queries. For example, for 1 million records, Q5 has approximately 100000 unique records of which the PRIs need to be computed; whereas for other queries, the algorithm only needs to compute PRIs for no more than a few thousand records and the rest can be directly retrieved using memoization.
5.3.2 Effect of parameter $p$
To find a suitable $\epsilon$, the algorithm computes the ratio of the maximum PRI over the PRI at the $p$th percentile; if the ratio is close to 1, the PRIs are considered to reflect adequate privacy protection by the controller’s privacy requirement. Specifically, we are interested in choosing the 00th (minimum), $25$th, and $50$th (medium) percentile. The smaller $p$ is, the stricter the privacy requirement and the harder it is to find $\epsilon$, meaning more noise needs to be added to achieve a ratio closer to 1.
Setup. We used the Adult dataset with 100000 rows that we generated in Section 5.3.1, and we used the same settings (except $p$). We ran the five queries in Section 5.3.1 10 times and we report the $\epsilon$ selected by our algorithm in each run.
Figure 10: $\epsilon$ found when using the PRI at the 0, 25, and 50th percentile to compare against the maximum PRI
Results. Figure 10 shows the $\epsilon$ found when $p=0,25,50$. As expected, using higher percentiles tends to find larger $\epsilon$ than lower percentiles, because less noise needs to be added to bring the maximum PRI closer to the PRI at the larger percentile. When using the minimum PRI, Q1 – Q4 find very small $\epsilon$ that are between 0 and 1. Q1 and Q3, both counting queries, are able to find larger $\epsilon$ since the queries have a global sensitivity of 1, i.e., the maximum difference between PRIs will also be fixed at 1, it becomes easier to make the ratio closer to 1. Whereas Q2 and Q4, a group-by query and an average query with higher sensitivities, require more noise to be added to bring the PRIs closer. For Q5, using the minimum PRI only finds $\epsilon$ once, since the sensitivity is too large, primarily due to the domain size of fnlwgt), the DP mechanism will not be able to add enough noise of such scale beyond certain $\epsilon$ (when $\epsilon=0.1$, the noise scale will be larger than 10,000,000, which is the threshold we use to determine whether the DP mechanism can be applied). However, when the requirement is relaxed, it is able to find larger $\epsilon$ such that the noise scale does not exceed the threshold.
6 Related Work
Choosing privacy loss budget. Different approaches [30] [26] [23] [36] [39] [40] [28] [48] [32] [28] aim to help users choose $\epsilon$ when using DP. One line of work [39] [40] [28] [48] [32] implements an interface to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. visualizes the disclosure risk in relation to $\epsilon$; they use the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ge et al. translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. proposes an economic model for choosing $\epsilon$ by balancing the interests of two different parties: the data analyst and the prospective participant who will choose to participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the "cost" of whether to participate in the study, which is impractical for non-experts. In addition, Cummings et al. and Xiong et al. investigate user’s understanding of DP guarantees through multiple user studies, which offer insights on facilitating communicating the privacy implications of DP to users and help them choose $\epsilon$. Jayaraman and Evans evaluates the impact of $\epsilon$ on ML tasks and specifically their privacy-utility tradeoffs. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
Interpreting utility. Tao et al. proposes DPXPlain to explain aggregate query answers under DP by comparing answers with and without DP, and assessing whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis. This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers. Moreover, feasibility studies [8][14] have been conducted to evaluate real-world usage of DP.
CDP LDP Shuffler
Crypto
Primitive
Data
Escrow
T?
No Yes Yes Yes
Yes
Add.
Trust
Asm.
Trust
indv’ls
add noise
correctly
Trust
shuffler
Trust
crypto
impl.
Trust
SEV &
escrow
impl.
Encr.
Data?
No
Does not
collect
raw data
Does not
collect
raw data
Yes Yes
Error $O(1)$ $O(\sqrt{n}$) Near $O(1)$ $O(1)$ $O(1)$
Table 1: Comparisons of different DP deployment scenarios. T stands for trustworthy.
Deploying differential privacy. Alternatives beyond the central model aim to deploy DP with fewer trust assumptions. We summarize those different deployment scenarios in Table 1. The local model (LDP) has been proposed to eliminate the need for a trusted data curator and it has been adopted in various real-world applications for collecting client data from their software and devices [24] [18] [17]. In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the data curator; the data curator does not need to be trusted since it only sees the differentially private version of data. However, the primary disadvantage of the local DP model is that it introduces a much larger error than the central DP model. The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees [34] [6], thus it requires less noise to be added when users perturb their data and offers better accuracy than the local DP model. However, it requires additional trust assumptions as the shuffler must be trusted. Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator [20] [42] [9] [4] [45] [46] [49] [10] [44] [11] [12]. We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
7 Conclusions
To ease the deployment of DP, we address challenges faced by controllers and analysts by 1) defining the privacy risk indicator (PRI) that indicates the relative privacy loss of each individual for a given $\epsilon$ and output and an algorithm to choose $\epsilon$ using the PRIs. 2) introducing an utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We propose a platform implementation on top of a data escrow. We demonstrated our contributions through an IRB-approved user study and empirical evaluation with other DP platforms.