### v2

# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

Abstract.
Differential privacy (DP) enables private data analysis but is hard to use in practice. For data controllers who decide what output to release, choosing the amount of noise to add to the output is a non-trivial task because of the difficulty of interpreting the privacy parameter $\epsilon$. For data analysts who submit queries, it is hard to understand the impact of the noise introduced by DP on their tasks.
To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm to choose $\epsilon$ and release output based on controllers’ privacy preferences; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows controllers to control dataflows while maintaining high performance. We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms. All in all, our work contributes to making DP easier to use by lowering adoption barriers.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a promising tool to analyze data about individuals, but it remains challenging to use in practice. In a typical DP deployment, a data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller. The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals. The data controller wants to protect individuals’ privacy while releasing useful output to the data analyst. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a few roadblocks prevent broader DP deployment. In this paper, we consider three such roadblocks:
Challenge 1: Choosing $\epsilon$. Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.  Dwork et al. showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
Our contribution. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects individuals’ privacy. And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release DP output given a simple description of controllers’ privacy preferences.
Challenge 2: Interpreting utility of DP output. Analysts submit queries and receive noisy output that deviates from the raw output, i.e., the ground truth output if DP had not been applied. Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks, which is the reason why they sought the DP output in the first place. This hinders the analysts’ ability to use the DP output in practice.
Our contribution. In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control data release. In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform. With the controller’s permission, the platform invokes the evaluation function on both the DP output and the raw output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
Challenge 3: Deploying our solutions in a DP platform. Deploying our solutions to the previous two challenges brings a new challenge on the platform design. Specifically, implementing the new Find-$\epsilon$-from-PRI algorithm (addressing Challenge 1) and utility signaling protocol (addressing Challenge 2) calls for a platform that meets the four requirements: (1) it allows controllers to control dataflows, (2) it implements necessary security measures to protect against unintended dataflows, (3) it facilitates controllers to inspect dataflows, and (4) it adds little overhead. Various DP platforms have been proposed, but none meet all four requirements. For instance, the central DP model assumes adversaries are incapable of corrupting the platform and forcing unintended dataflows, which is unrealistic in practice (Near et al., 2021); platforms leveraging cryptographic primitives (Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018) make fewer assumptions about the adversaries, but often have significant computation overhead.
Our contribution. We propose a new platform built on top of a data escrow (Xia et al., 2022) which enables the implementation of the Find-$\epsilon$-from-PRI algorithm and utility signaling protocol, meeting the four platform requirements we curated. We show that the escrow-based platform hits a new point in the design space, offering a better trade-off between security, usability, and performance.
[…]
Summary of Contributions and Related Work
Helping controller choose $\epsilon$ using PRI. Many approaches have been proposed to help controllers choose $\epsilon$ (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Ligett et al., 2017), however, none concentrate on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do. In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users who share their data (Nanayakkara et al., 2023; Franzen et al., 2022), whereas our approach focuses on helping controllers, who can only make an informed choice of $\epsilon$ if they understand its privacy implications on those end-users.
Helping analyst via utility signaling protocol. Adding DP-style noise to the output affects the quality and usability of the data for the analyst’s downstream tasks (Hotz et al., 2022). For instance, the adoption of DP by the US Census has raised major concerns regarding the reliability of the census statistics after noise infusion and its effect on research and policy-making (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]). Our utility signaling protocol takes a step forward in helping analysts interpret the effect of DP on the released output.
Escrow-based DP platform. Various DP platforms have been proposed (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017), but none meets the requirements for deploying our solutions to help controllers and analysts. Our escrow-based platform meets those requirements, enabling practical deployment of DP.
Outline. The rest of the paper is organized as follows. We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3. Section 4 presents the utility signaling protocol and Section 5 presents the escrow-based platform. We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
2. Privacy Risk Indicator (PRI)
Background. Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual. Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O).$
One way of achieving differential privacy is by constructing a mechanism that adds noise to the query output. The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff (Near et al., 2021).
2.1. Why Choosing $\epsilon$ is Hard and PRI Overview
Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy. The parameter $\epsilon$ bounds the maximum privacy loss for any individual, but it is challenging for controllers to reason about the actual privacy implications of releasing an output for specific individuals in the dataset after choosing $\epsilon$ (Dwork et al., 2019).
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output. We construct PRI so that it is directly connected with the concept of privacy loss, which we explain next. This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$. One strategy includes an algorithm to find $\epsilon$ based on the controller’s privacy preferences expressed in terms of PRIs, which we will explain in Section 3.
2.2. Privacy Risk Indicator: Formal Definition
We represent a dataset as a set of $n$ records $x=(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, drawn from a data universe $\mathcal{X}$. WLOG, we assume each individual (labeled $1$ to $n$) contributes exactly one record to $x$.
We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
Definition 0 (Privacy Risk Indicator).
The privacy risk indicator of an individual $i$ is defined as:
$PRI_{i}=\lVert r-q(x_{-i})\rVert_{1},$
where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$.
Connection between PRI and privacy loss. Next, we show that when the Laplace Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of individual privacy loss. First, we show the definition of privacy loss in DP (Dwork et al., 2014).
Definition 0 (Privacy Loss).
Given two neighboring datasets $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
$\mathcal{L}_{\mathcal{M}(x,q,\epsilon)\|\mathcal{M}(x^{\prime},q,\epsilon)}^{(r)}=\ln\left(\frac{\Pr\left[\mathcal{M}(x,q,\epsilon)=r\right]}{\Pr\left[\mathcal{M}(x^{\prime},q,\epsilon)=r)\right]}\right).$
Next, we introduce the definition of Laplace Mechanism (Dwork et al., 2014), an $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
Definition 0 (Laplace Mechanism).
Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon>0$, the Laplace Mechanism is defined as:
$\mathcal{M}_{L}(x,q,\epsilon)=q(x)+(Z_{1},\ldots,Z_{k}),$
where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{GS_{q}}{\epsilon}$  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
Using the Laplace Mechanism, we express the relative privacy loss of individual $i$ as:
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}$ $=\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/GS_{q})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/GS_{q})}\right)$
$=\frac{\epsilon(\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1})}{GS_{q}}.$
Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}>\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-j},q,\epsilon)}^{(r)}.$
The proof is straightforward: if $PRI_{i}>PRI_{j}$, then $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$, thus the inequality holds. Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss of any individual will always be bounded by $\epsilon$.
We now make several remarks:
PRI gives controllers additional information to decide what to release. The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output. In this paper, we show how PRI helps controllers decide which output to release.
PRI is data-dependent and thus leaks information about individuals. Controllers already have access to individuals’ raw data, so when they gain access to the PRI, they do not learn anything extra about the individuals. But this means the PRIs must remain private to the controllers and never be shared with anyone else.
Controllers may give PRI different uses. Controllers may express their privacy preferences for individuals in the dataset in terms of their associated PRIs, and derive an $\epsilon$ that produces that set of PRIs. This means that eliciting controllers’ privacy preferences based on PRIs paves the way for choosing $\epsilon$. We exploit this insight in the next section to design an algorithm to find $\epsilon$. We also discuss other ways in which PRIs may be used in Section 3.5.
3. Deriving Epsilon based on the PRI
In this section, we describe an algorithm to choose $\epsilon$ given controllers’ privacy preferences in terms of PRIs, thus addressing Challenge 1 (Sections 3.1, 3.2). In Section 3.3, we discuss how to bound the total privacy loss in the presence of multiple queries. In Section 3.4, we provide the end-to-end DP guarantee for the algorithm. We discuss the algorithm’s complexity and extensions in Section 3.5.
3.1. Using PRI to Express Privacy Preferences
Traditionally, controllers need to choose $\epsilon$ to convey their privacy preferences. However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset. In contrast, the per-individual PRI indicates the relative privacy loss of an individual compared to other individuals in the dataset. The consequence is the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output. Next, we convey the intuition on how the PRI is useful to controllers through a vignette.
Figure 1. Example of using PRI to find $\epsilon$. For illustrative purposes, we show the PRI of each student (with the corresponding query output) and leave out students’ actual GPA.
Vignette illustrating one use of the PRI. Figure 1 shows the resulting PRIs of four students after answering a query about their average GPA. If the query output were released without using DP (a), the analyst observing the output would learn a lot about those students’ grades, as indicated by the PRI (remember that the PRI is a proxy to privacy loss). If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism. This would lead to a noisy output and a PRI associated with each student (note that running the mechanism two times with the same $\epsilon$ does not guarantee the same PRIs are obtained). We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students. In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others. Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.2) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve. Lastly, the PRIs in both (c) and (d) indicate no student will be at a higher risk than others, but (d) produces a query output with worse accuracy. The controller may choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
The above is an illustrative example: finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
3.2. The Find-$\epsilon$-from-PRI Algorithm
When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the relative privacy loss of each individual compared to others. Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a set of PRIs that represents their preferences. We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI. In Section 3.1, we have shown one criterion for choosing $\epsilon$ that leads to exactly equal PRIs, which is typically infeasible. Instead, we propose an alternative criterion to let controllers decide whether a query output respects their privacy preferences. We explain the key idea and then the algorithm.
Key Idea. We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset. This is done by requesting exactly one parameter from the controllers, a percentage p ($0\leq p\leq 100$). We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$. Then we find an $\epsilon$ that produces a list of PRIs (and associated DP output) such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$. In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI. For instance, setting $p=0$ means the PRIs should be exactly equal.
Unlike $\epsilon$, the parameter $p$ directly conveys the level of risk conferred to individuals after releasing a specific query output. Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$. The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
Output : Chosen privacy parameter $\epsilon$, query output $r$
1 /* Initialize set of $\epsilon$ values to test */
2 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
3 /* Compute $q(x_{-i})$ for all individual $i$ using memoization */
4 $C\leftarrow map()$
5 for $x_{i}\in x$ do
6       if $x_{i}\notin C$ then
7             $C[x_{i}]\leftarrow q(x_{-i})$
8
9 for $\epsilon\in\mathcal{E}$ do
10       /* Compute query output using DP with current $\epsilon$/
11       $r\leftarrow\mathcal{M}_{L}(x,q,\epsilon)$
12       /* Compute corresponding PRI of each individual */
13       $PRIs\leftarrow list()$
14       for $x_{i}\in C$ do
15             $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
16             $PRIs.add(PRI_{i})$
17      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
18       /* Check whether $\epsilon$ satisfies controller’s privacy preference */
19       if $PRI_{min}/PRI_{max}\geq 1-(p/100)$ then
20             return $r$
21
Algorithm 1 Find-$\epsilon$-from-PRI
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ values is infinite, we set an upper bound and a range, e.g., 10 and 0.1, to ensure a finite set of options, e.g., 100 in this example. Then, for each $x$’s neighboring dataset $x_{-i}$, the algorithm computes $q(x_{-i})$ (line 1 - 1), which will be used to compute the PRI for each individual $i$. Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1). For each $\epsilon$, the algorithm computes the query output $r$ by applying the Laplace Mechanism $\mathcal{M}_{L}$ with parameter $\epsilon$ (line 1) and computes the corresponding PRI of each individual (line 1 - 1). Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1). If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output.
Figure 2. Example of $\epsilon$ chosen by the Find-$\epsilon$-from-PRI algorithm with different percentage $p$ over 50 runs
Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRI profiles across executions of the mechanism; simultaneously, similar PRI profiles may be produced by different $\epsilon$ values. What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$. In addition, we show empirically in Section 6.3.2 that stricter privacy preferences (i.e., smaller $p$) generally lead to smaller $\epsilon$. Figure 2 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs; for demonstration, we choose one of the five queries (Q5) we have run. As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$. Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
Figure 3. Dataflow of Find-$\epsilon$-from-PRI Algorithm
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm. Individuals send their data to the controller; this dataflow remains unchanged from traditional central DP models (Near et al., 2021). The analyst submits a query to the controller. The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to find a suitable query output, which is an output produced by a standard DP mechanism (Laplace Mechanism in our work) with a value for $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The PRI associated with each individual is never released to the analyst. The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval. The value for $\epsilon$ that led to the query output is also not released by default.
Releasing $\epsilon$. Under certain scenarios (Dwork et al., 2019), $\epsilon$ may also need to be released. Since $\epsilon$ is chosen based on PRIs, which are data-dependent, the controller should release $\epsilon$ differential privately by incorporating techniques such as Exponential Mechanism (e.g., Report Noisy Max) (McSherry and Talwar, 2007) or Sparse Vector Technique (Dwork et al., 2009; Lyu et al., 2016) in the algorithm. Instead of choosing the largest $\epsilon$ that passes the PRI criterion, the algorithm will choose and return the first $\epsilon$ (which may or may not be the largest) that passes the PRI criterion while preserving DP.
3.3. Bounding Privacy Loss Across Queries
In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery (Dwork et al., 2014). In today’s deployment of DP, controllers need to set a total privacy budget on the dataset a priori and stop accepting queries when the cumulative $\epsilon$ exceeds the budget (Dwork et al., 2019). The need for setting a total budget poses a critical challenge for controllers:
Controllers do not know the privacy implications of total privacy budget. Similar to choosing $\epsilon$ for a single query, when deciding which total budget to set, controllers face the same challenge of not knowing the actual privacy implications on individuals after releasing multiple query outputs. Oftentimes, they resort to selecting the total budget arbitrarily or through trial-and-error (Dwork et al., 2019).
The solution. Instead of quantifying the maximum acceptable privacy loss, which is hard, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$. For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). The algorithm rejects the query if it fails to find a suitable $\epsilon$ larger than $\epsilon_{c}$, meaning the controller’s privacy preferences cannot be met for any possible $\epsilon$.
Our solution creates a dynamic bound of total privacy loss. As $\epsilon_{c}$ keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since larger $\epsilon$ may not produce PRIs that satisfy controllers’ privacy preferences. At each point the algorithm fails to find a suitable $\epsilon$ for a query, the total privacy loss is bounded. The bound is refreshed if a suitable $\epsilon$ is found for a different query. A potential concern is whether there is additional privacy leakage using this dynamic bound. In Section 3.4, we show the bound can be formalized as a privacy odometer in fully-adaptive composition (Rogers et al., 2016) and there is no leakage.
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output with degrading accuracy. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ in the algorithm based on analysts’ preferences, e.g., they can set a higher $p$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
3.4. End-to-End DP Guarantee
Implications of choosing $\epsilon$ adaptively. First of all, if the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ that is chosen by the algorithm. However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$. Then the total privacy loss is not bounded by any fixed constant since the algorithm does not require setting a total privacy budget. This brings the question of whether releasing multiple query outputs computed using adaptively chosen $\epsilon$ causes additional privacy leakage.
Privacy odometer in fully-adaptive composition. To answer this question, we construct a privacy odometer (Rogers et al., 2016) for the algorithm, which provides a probabilistic running upper bound on the realized privacy loss after answering each query. The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
Theorem 3.4 in  (Rogers et al., 2016) has proved that there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition. Therefore, the privacy odometer for the Find-$\epsilon$-from-PRI algorithm under sequential composition (Dwork et al., 2014) can simply be defined as:
Definition 0 (Privacy Odometer).
[…]
The privacy odometer gives the same privacy bound as sequential composition; since the algorithm uses $\epsilon$-DP mechanism, $\delta_{g}$ (the global failure parameter in $(\epsilon,\delta)$-differential privacy) can be set to 0. The privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}$ after the algorithm answers each query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $\epsilon_{c}$-differential privacy.
It is worth noting that although  (Rogers et al., 2016) showed the privacy odometer gives a worse bound on privacy loss than the standard advanced composition theorem (Dwork et al., 2010), more recently (Whitehouse et al., 2023) proved that in the fully-adaptive setting the privacy odometer can achieve the same asymptotic bound as advanced composition under certain conditions. For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition.
3.5. Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel. Thus, we implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 6.3.1 that the algorithm is both efficient and scalable.
Expressing privacy preferences. By default, the algorithm considers the list of PRIs for all individuals in the dataset when finding $\epsilon$. A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of records. For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x^{\prime}$) and those younger ($x^{\prime\prime}$), and specify that the two groups need similar privacy protection. This preference can easily be incorporated into the algorithm by letting the controller change the two PRI values to compare instead of the global $PRI_{min}$ and $PRI_{min}$ for all individuals; e.g., they can compute $PRI_{med}^{\prime}$ and $PRI_{med}^{\prime\prime}$, the median PRI for individuals in $x^{\prime}$ and $x^{\prime}$ respectively, and choose an $\epsilon$ that makes $PRI_{med}^{\prime}/PRI_{med}^{\prime\prime}\geq 1-(p/100)$.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP. The PRI complements the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
4. Helping Analysts Interpret Utility
In this section, we concentrate on solutions to help analysts interpret the utility of DP output. We revisit Challenge 2 (Section 4.1) and present a solution we call utility signaling protocol (Section 4.2).
4.1. Challenge Overview and Key Insight
Controllers work on a platform that has access to the individuals’ data and that is responsible for answering analysts’ queries using DP. An analyst submits a set of queries $Q=\{q\mid q\in\mathcal{Q}\}$ ($\mathcal{Q}$ is the query space) to the platform, and (if a suitable $\epsilon$ is found for each query) receives a set of noisy query output $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n},q\in Q\}$ ($\mathcal{M}$ is a DP mechanism and $\mathcal{X}$ is the data universe). A well-known challenge of using DP is that analysts do not know the utility of $S$ (Hotz et al., 2022). Concretely, analysts do not know the effect of using $S$ on their downstream task, which is the purpose for analysts to query the data in the first place.
Example 1/3. An analyst queries census data at the county level. The set of query output $S$ is aggregate data on population count, education level, and housing prices that are released using DP. The analyst uses $S$ to train a binary classifier to predict the income level of each county and obtains an accuracy of 68%. The analyst does not know if the (low) accuracy is due to the effect of DP or inappropriate training data.
The difficulty in interpreting the utility of $S$ makes it challenging for analysts to understand how the DP output will impact their downstream tasks; this is one of the core issues surrounding the discussion of the use of DP for the US Census (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
Observation. Analysts use the DP output $S$ for some downstream task, such as a machine learning or inference task. Downstream tasks often have an associated evaluation function, $g(\cdot)\rightarrow\mathbb{R}$, which returns a number indicating the quality of the downstream task when evaluated on a certain input. In the example above, the downstream task is to train a binary classifier, the evaluation function measures the quality of the classifier trained on DP output and returns 68% as the quality indication (accuracy in this case).
Insight: signaling DP effect. Analysts would learn the impact of DP if they had access to both the result of the evaluation function on the DP output, $g(S)$ [^2] We slightly abuse notation for simplification. and the raw output, $g(S^{\prime})$, where $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n},q\in Q\}$ (the set of query output without using DP).
Example 2/3. Besides having access to the quality of the classifier on the DP census aggregates (68%), the analyst also gains access to the quality of the classifier if it had been trained on the raw census aggregates $S^{\prime}$, say 80%. By comparing 68% ($g(S)$) and 80% ($g(S^{\prime})$), the analyst learns the effect of added DP noise on their downstream task.
Gaining access to $g(S)$ and $g(S^{\prime})$ helps analysts interpret the impact of DP on their downstream tasks. The remaining challenge is how to provide $g(S^{\prime})$ to the analysts without giving them access to $S^{\prime}$, which is the raw query output and cannot be released to the analyst without violating the privacy of individuals in the dataset.
4.2. The utility signaling protocol
Rather than the platform directly sending $S^{\prime}$ to the analyst, which would break individuals’ privacy, the analyst submits their downstream task and associated evaluation function $g(\cdot)$ to the platform. Upon successful inspection by the controller, the platform executes the downstream task and evaluation function on behalf of the analyst, producing $g(S)$ and $g(S^{\prime})$. Last, if the controller approves, the platform releases both $g(S)$ and $g(S^{\prime})$ to the analyst, thus signaling the impact of DP on the analysts’ task without revealing $S^{\prime}$.
Example 3/3. Instead of only submitting the queries to the platform, with the utility signaling protocol, the analyst submits their queries, the downstream task, and the evaluation function. The controller inspects each and allows the platform to signal the impact of using DP to the analyst by releasing both utility results 68% and 80%.
4.2.1. Protocol Step by Step
Figure 4. End-to-end example of utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
Figure 4 illustrates the utility signal protocol, which we describe next:
0. Platform. The platform has access to the dataset $x$ of individuals. It also has the capacity to run queries using DP. It may optionally implement the Find-$\epsilon$-from-PRI algorithm presented in Section 3.2333While we envision the algorithm being deployed to help controllers, both contributions are orthogonal.
1. Analyst submits queries to the platform. The platform notifies the controller.
2. Controller inspects queries. The controller inspects the analyst’s queries and allows the platform to execute the queries as long as a suitable $\epsilon$ is selected. The platform withholds both the set of DP output $S$ and raw output $S^{\prime}$ without releasing anything to the analyst by default.
3. Platform releases $S$ upon controller’s approval. The platform notifies the controller after computing $S$ and $S^{\prime}$. The controller gives permission for the platform to release $S$ to the analyst.
Up to this point, the protocol provides the analyst with the same information they can access in a typical central DP deployment. Now, we extend the protocol to implement the utility signaling component to help the analyst understand the impact of DP on their downstream task.
4. Analyst submits downstream task and evaluation function $g(\cdot)$ to the platform. The platform notifies the controller.
5. Controller inspects downstream task and evaluation function. The controller inspects the implementation of the analyst’s downstream task and evaluation function $g(\cdot)$, and decides whether it is appropriate to release $g(S)$ and $g(S^{\prime})$. With the controller’s approval, the platform executes the downstream task and evaluation function $g(\cdot)$ on both $S$ and $S^{\prime}$, withholding the results by default.
6. Platform releases $g(S)$ and $g(S^{\prime})$ upon controller’s approval. The controller can inspect $g(S^{\prime})$ one last time to ensure it is appropriate to release; $g(S)$ preserves DP since it is a post-processing (Dwork et al., 2014) step of DP output $S$. If approved, the platform releases $g(S)$ and $g(S^{\prime})$ to the analyst, who learns the impact of DP on their downstream task without ever gaining access to $S^{\prime}$ directly.
4.2.2. Privacy implications of releasing $g(S^{\prime})$
Any data release leaks information. When releasing an output that is subject to DP, such as $g(S)$, controllers understand the consequences on the individual’s privacy. Conversely, when releasing output not subject to DP, such as $g(S^{\prime})$, controllers must assess whether doing so maintains individuals’ privacy. That ultimately depends on the dataset $x$, the output $S^{\prime}$, the implementation of the downstream task and evaluation function, $g(\cdot)$, and the adversary model. The scope of the utility signaling protocol is to enable analysts to understand the impact of using DP, but not on understanding, controlling, and determining when it is safe to release $g(S^{\prime})$, which is ultimately at the discretion of the controllers and an interesting avenue of future work.
5. Escrow-Based DP Platform
We curate a set of requirements for a DP platform that implements the Find-$\epsilon$-from-PRI algorithm and the utility signaling protocol in Section 5.1. In Section 5.2, we describe the implementation of a platform built on top of a data escrow that meets these requirements.
5.1. Platform Requirements
The solutions to Challenge 1 (the Find-$\epsilon$-from-PRI algorithm), to Challenge 2 (the utility signaling protocol), and the nature of processing over private data call for a platform that meets the following requirements:
  • R1. Control dataflows: Given the three relevant entities (analyst, platform, and controller), a dataflow indicates data movement between any two of those entities. A requirement for the platform is to ensure that only permitted dataflows take place. Concretely, controllers indicate what dataflows take place, and the platform enforces the controller’s decision. In the context of our work, analysts can access for example the DP output $S$ after the controller approves, and they cannot access PRIs or any other intermediates not explicitly approved by the controller.
  • R2. Secure platform: Adversarial agents may break the requirement of controlling dataflows by breaking into the platform and forcing unintended dataflows. This is an assumed risk in the central DP model, which can be reduced by using a local DP deployment (Near et al., 2021). Unfortunately, local DP reduces the accuracy of the output compared to central DP. A requirement for the platform is to ameliorate the risk of unintended dataflows.
  • R3. Protect against adversarial downstream and evaluation functions: Adversarial analysts could submit downstream tasks and evaluation functions crafted to release more information than needed. While controllers must inspect these functions before agreeing to release the output and they are the ultimate arbiter, a requirement of the platform is to facilitate the controllers’ job.
  • R4. High Performance: The Find-$\epsilon$-from-PRI algorithm increases processing overhead. Simultaneously, security safeguards to address R2 are often at odds with high-performance computation. Together, these two facts risk making the platform impractical. A requirement is to implement the solutions to Challenge 1, Challenge 2, and meet the requirements R1, R2, and R3 while maintaining low overhead.
We now explain the implementation of a platform that addresses these three requirements.
5.2. Platform Design and Implementation
To demonstrate the feasibility of the solutions to Challenge 1 and Challenge 2, we describe how to implement the solutions above on a platform that leverages a data escrow (Xia et al., 2022) architecture.
Data escrow overview. A data escrow is a system that acts as an intermediary for storing, processing, and releasing data. Different entities interact within the escrow. Controllers register data and control the dataflows. The data is never stored or computed in the clear using end-to-end encryption. Analysts run computations on registered data in the escrow. For any data or its derived data product to leave the escrow, controllers have to create a policy on the escrow. The policies are tamper-proof since the escrow validates the identity of the controller who creates the policy and the policy remains encrypted, thus ensuring the controller fully controls the dataflow.
Meeting R1. No raw data (e.g., $x$) or derived data product (e.g., PRIs, $g(S)$, $g(S^{\prime})$) leave the escrow without an explicit policy that permits that. Policies then determine what dataflows take place. Controllers are in charge of writing policies.
Meeting R2. Data is encrypted end-to-end, even during computation. Adversaries that break into the platform cannot read plain-text data. In other words, no unintended dataflows take place.
Meeting R3. The platform can let analysts submit their downstream task and associated evaluation function, and put the burden on the controller to verify the implementation of those functions. Or it can offer standard implementation of downstream tasks and evaluation functions and let analysts call these, thus requiring scrutinizing the functions once and using them many times. The data escrow permits the deployment of programs expressing a desired functionality. In specific DP scenarios, e.g., the US Census, it is conceivable that one could provide programs implementing (at least) the more standard functionalities and query interfaces, and rely on controllers to scrutinize only one-off tasks.
Meeting R4. End-to-end data encryption (necessary to meet R2) can be achieved in many ways; for instance, DP platforms leveraging Secure Multiparty Computation (Narayan and Haeberlen, 2012; Bater et al., 2016, 2018; Wang et al., 2022; Roth et al., 2019), Homomorphic Encryption (Agarwal et al., 2018; Roy Chowdhury et al., 2020), Oblivious RAM (Bogatov et al., 2021) and Trusted Execution Environment (Nguyen et al., 2020). Our evaluation prototype uses AMD Secure Encrypted Virtualization (Kaplan et al., 2016) because it permits running unmodified software while maintaining the integrity and confidentiality of memory contents. The escrow design protects access to external resources (e.g., network and disk), ensuring data is encrypted transparently to any application. As we show in the evaluation (Section 6.2), this solution virtually removes the processing overhead. In addition, we implement a series of optimizations to the Find-$\epsilon$-from-PRI algorithm to demonstrate its efficiency and scalability when run in the escrow (Section 6.3.1).
6. Evaluation
Our evaluation shows our solutions address Challenge 1 to help controllers, and Challenge 2 and 3 using an escrow-based platform. We answer the following research questions:
  • RQ1: Does presenting the PRI help data controllers choose $\epsilon$?
[…]
  • Microbenchmarks: We conduct experiments to show that
    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study the effect of varying parameter $p$ in the algorithm.
6.1. RQ1: Using PRI To Choose Epsilon
To determine whether PRI helps the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$. To this end, we conducted a between-subjects user study where the participants are divided into control and treatment groups. To prevent learning effects, only the treatment group sees the PRI. We chose to conduct the study in the form of an online survey because our tasks do not require in-person guidance and it is more convenient for participants to complete the study. In addition, it allows us to collect data from a more representative sample.
6.1.1. Recruitment and data collection
We recruited 100 participants using Prolific (pro, [n.d.]). To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP address, location data, and contact information are not being recorded.
Prescreening criteria. When recruiting on Prolific, we prescreen participants based on three criteria: i) they are fluent in English, to avoid confusion due to misunderstanding language; ii) they have completed an undergraduate degree or above; iii) they study Computer Science, Computing (IT) or Mathematics; this is to ensure that the participants have a sufficient quantitative background to complete the study.
Tutorial and Screening Quiz. Since the main goal of the study is to ask the participant to choose $\epsilon$ when using DP to release query output, the participant needs to have basic knowledge about DP and understand the implications of the $\epsilon$ they chose. To ensure each participant knows enough about DP to complete the study, we ask each participant to read a short tutorial about DP at the beginning. This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level. Mainly, it explains that DP is generally achieved by adding random noise to the query output, and larger $\epsilon$ means less noise will be added and vice versa.
After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read. Only participants who pass the quiz proceed to conduct the study. Of the 100 participants we recruited, 56 passed the quiz and their responses were used for the study analysis. This quiz ensures that participants have understood the principles of DP, thus strengthening the validity of the user study.
6.1.2. Study procedure
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group. The survey presents a sample from the UCI Adult dataset (uci, [n.d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order:
  • The number of people with income greater than 50K and are less than 40 years old.
  • The number of people who have a Master’s degree or above and have income $\leq$ 50k, grouped by each race.
For each query, the survey asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 in an interval of 0.5. Since we want all participants to have a clear task goal in mind, the survey includes a description of the task: choose $\epsilon$ to protect individuals’ privacy while yielding as accurate answers as possible.
For the control group, the participant is presented with a table showing the range of possible output for each $\epsilon$ to indicate output accuracy and the raw (non-DP) output. The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
For the treatment group, the participant is presented with the same table, but with an additional column showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI); we also provide a brief explanation of the PRI. Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
6.1.3. Results
Figure 5. $\epsilon$ chosen by each participant
If the intervention (showing the PRIs) is effective, we expect this would manifest as more consistent answers among participants, since all participants were given the same training and the same task. Therefore, we compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups.
Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control or treatment group. For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group. This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy. This shows that most participants could connect the PRIs and their task goal, which is to protect individuals and release useful output, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal. On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to choose.
We ran Mann–Whitney test  (McKnight and Najab, 2010) on the $\epsilon$ values chosen from the control and treatment group, with the alternative hypothesis being the two samples come from different distributions. For both queries, we obtained $p<0.005$. Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
Last, we asked participants in the treatment group if they used the PRIs to complete the task: all 27 participants responded they did use that information, confirming they were indeed “treated”.
Figure 6. $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
Does knowing DP matter? Another interesting question is whether the participant’s prior familiarity with DP affects how they choose $\epsilon$. At the start of the survey before showing the tutorial, we asked each participant whether they know about DP. 27 participants (out of a total of 56) answered they do not know anything about DP and 29 participants answered they know about DP.
Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity. We find similar patterns. Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance. This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they know about DP previously.
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$ instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
Figure 7. $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference. Participants who saw histograms did not perform better than participants who saw range plots. Especially for Q2, eight participants chose larger $\epsilon$ (3.5, 4, and 4.5). Since the responses are anonymous, we cannot directly ask participants about their decision-making after the survey ends. We deduce one possible explanation might be that for the large $\epsilon$, the majority of PRIs are already small and similar, and only a small portion of PRIs have large values. The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate. However, if the participants only saw the range of PRI, they would be more inclined to choose smaller $\epsilon$ with a smaller range of PRI. Another possible explanation could be that some participants had difficulties interpreting histograms, whereas the range plots are generally much easier to understand.
The discrepancy of results by tweaking how we present the PRIs demonstrates how we convey information to the users affects their ability and reasoning for choosing $\epsilon$. Exploring other alternative solutions for presenting and helping users interpret PRI remains an interesting problem.
6.1.4. Threats to validity
External validity: We recruited participants from diverse backgrounds worldwide via Prolific, with few prescreen criteria so that the participants could complete the study without difficulty. Internal validity: To ensure participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz. Participants can only take the quiz once to minimize the chance of them abusing it if they have multiple chances. To avoid learning effects, we used a between-subject study to isolate the condition we intended to test. We also randomized the two questions for choosing $\epsilon$.
6.2. RQ2: DP On Escrow-based Platform
[…]
6.2.1. Baselines for comparison
We evaluate and compare the performance of the four following baselines:
  • Data Escrow: We implemented an escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output. Optionally, if the parameter $\epsilon$ is not given, it will use the Find-$\epsilon$-from-PRI algorithm to select a suitable $\epsilon$. We simulate the two entities (controller and analyst) in the escrow. A policy is first created by the controller to allow the query output to be released to the analyst.
[…]
  • Standalone with DP: This baseline runs the given query and applies the Laplace Mechanism to the query output in a central platform. The dataset is in plaintext. This represents the central DP deployment scenario without the overhead of additional security measures.
  • Standalone without DP: This baseline simply runs the given query without DP in a central platform. The dataset is in plaintext. We include this baseline to show the overhead of using DP.
[…]
6.2.2. Setup
We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine. We used AMD Secure Encrypted Virtualization (SEV) when running the data escrow. We used Ubuntu 20.04 and Python 3.8 for all baselines.
6.2.3. Dataset
[…]
6.2.4. Queries
We run the 3 queries independently for all baselines:
  • Q1: Number of people between age 20 and 30
  • Q2: Number of people by each race
  • Q3: Number of females who have income greater than 50K
6.2.5. Results
[…]
Figure 8. Average query time (in log scale) over 10 runs
[…]
6.3. Microbenchmarks
We evaluate the runtime and scalability of the Find-$\epsilon$-from-PRI algorithm in Section 6.3.1 and study how different values of the parameter $p$ affect the $\epsilon$ chosen in Section 6.3.2.
6.3.1. Is Find-$\epsilon$-from-PRI practical?
Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical. We ran the algorithm in the data escrow to further demonstrate its performance for computation-heavy tasks.
Datasets. We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset. We sampled records for datasets of sizes 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all attributes in the dataset.
Queries. We ran the algorithm to find $\epsilon$ for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.
  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num==13 AND age==25
  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==
    ’Asian-Pac-Islander’ AND 30<=age<=40 GROUP BY marital_status
  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=
    ’United-States’ AND sex==’Female’
  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass in (’Federal-gov’, ’Local-gov’, ’State-gov’)
  • Q5: SELECT SUM(fnlwgt) FROM adult WHERE capital_gain>0 AND income==’<=50K’ AND occupation==’Sales’
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total). We set $p=50$. We used six parallel processes for computing the PRIs.
Figure 9. Average runtime (in log scale) to find $\epsilon$ over 10 runs
Results. We first note that all queries on all 10 runs were able to find a suitable $\epsilon$ on every dataset we tested. Figure 9 shows a line plot where the x-axis is the data size and the y-axis is the runtime, both in log scale. Each line represents a query, with the error band showing the standard deviation.
As shown in the figure, the runtime increases linearly with the data size. Most queries took less than a few minutes to finish, even for 1 million records, except Q5, which took more than an hour. The runtime for Q5 is slower since it involves the attribute fnlwgt, which is a numerical attribute with a large domain size and many unique values, thus the query involves significantly more unique records than the other queries. For example, for 1 million records, Q5 has approximately 100000 unique records of which the PRIs need to be computed; whereas for other queries, the algorithm only needs to compute PRIs for no more than a few thousand records and the rest can be directly retrieved using memoization.
6.3.2. Effect of parameter $p$
The Find-$\epsilon$-from-PRI algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$. Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer. We test the effect of varying $p$ in the algorithm; specifically, when $p=5,25,50,75,95$.
Setup. We use the Adult dataset with 100000 rows that we generated in Section 6.3.1. We ran each of the five queries from Section 6.3.1 for 50 times, and report the $\epsilon$ chosen by our algorithm in each run.
Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$
Results. Figure 10 shows boxplots of $\epsilon$ found by the algorithm for different $p$. As shown in the figure, setting larger $p$ tends to find larger $\epsilon$. Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 for Q1–4 and less than 1 for Q5. Queries with larger sensitivity (Q2, Q4, Q5) can find larger $\epsilon$ when $p>50$, and queries with sensitivity of 1 (Q1, Q3) tend to find smaller $\epsilon$. Overall, the results empirically demonstrate that the algorithm is useful for finding a suitable $\epsilon$ (and corresponding query output) that is consistent with the controller’s privacy preferences.
7. Related Work
Defining PRI. The PRI shares similar intuition as Personalized Differential Privacy (PDP) (Jorgensen et al., 2015) in that different individuals may need different levels of protection since some records are more sensitive than others. In contrast to PDP, which requires each individual to specify their own privacy requirement through a privacy budget, the PRI is computed by controllers to indicate the realized privacy loss of each individual if a DP output were to be released.
Choosing privacy loss budget. Different approaches (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Hay et al., 2016; Ligett et al., 2017) aim to help users choose $\epsilon$ when using DP. One line of work (Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021) implements interfaces to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ligett et al. empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting. Ge et al. translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts. Cummings et al. and Xiong et al. investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
Interpreting utility. Tao et al. proposed DPXPlain to explain aggregate query answers by comparing answers with and without DP and assess whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis. This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers.
CDP LDP Shuffler
Crypto
Escrow
T
No Yes Yes Yes
Yes
Add.
Trust
Asm.
Trust
individuals
add noise
correctly
Trust
shuffler
Trust
crypto
impl.
Trust
SEV &
escrow
impl.
Encr.
Data?
No
Does not
collect
raw data
Does not
collect
raw data
Yes Yes
Error $O(1)$ $O(\sqrt{n}$) Near $O(1)$ $O(1)$ $O(1)$
Table 1. Comparisons of different DP deployment scenarios. T stands for trustworthy.
Deploying differential privacy. The threat model of the central DP model assumes adversaries cannot breach the platform and force unintended data flows, which is unrealistic. Alternatives beyond the central model aim to deploy DP with different trust assumptions. We summarize those different deployment scenarios in Table 1. The local model (LDP) has been proposed to eliminate the need for a trusted platform and it has been adopted in various real-world applications for collecting client data from their software and devices  (Erlingsson et al., 2014; Ding et al., 2017; Differential Privacy Team, 2017). In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the platform; adversaries can control the platform but they only gain access to DP data. However, LDP also introduces a much larger error ($O(\sqrt{n})$, where n is the number of individuals) than the central DP model ($O(1)$). The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees  (Kasiviswanathan et al., 2011; Balle et al., 2018), thus it requires adding less noise when users perturb their data and it offers better accuracy than the local DP model. However, it requires additional trust assumptions as the shuffler must be trusted. Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator  (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017; Bogatov et al., 2021). We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
8. Conclusions
In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain additional understanding of the effects of applying DP to the results of an analyst’s query. This translates into less burden choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI.
In addition, we show that it is possible to augment analysts’ understanding of the effects of DP on their results by using a utility signaling protocol that reveals the isolated impact of DP on a downstream task. We then gather the requirements gleaned from the previous two contributions and propose a platform, based on a data escrow, to deploy both ideas.