### v3

# Making Differential Privacy Easier to Use for Data Controllers using a Privacy Risk Indicator

Abstract.
Differential privacy (DP) enables private data analysis but is difficult to use in practice. In a typical DP deployment, data controllers manage individuals’ sensitive data and are responsible for answering data analysts’ queries while protecting individuals’ privacy; they do so by choosing $\epsilon$, the privacy loss budget, which controls how much noise to add to the query output. However, it is challenging for data controllers to choose $\epsilon$ because of the difficulty of interpreting the privacy implications of such a choice on the individuals they wish to protect.
To address this challenge, we first derive a privacy risk indicator (PRI) directly from the definition of ex-post per-instance privacy loss in the DP literature. The PRI indicates the impact of choosing $\epsilon$ on individuals’ privacy. We then leverage the PRI to design an algorithm to choose $\epsilon$ and release query output based on data controllers’ privacy preferences. We design a modification of the algorithm that allows releasing both the query output and $\epsilon$ while satisfying differential privacy, and we propose a solution that bounds the total privacy loss when using the algorithm to answer multiple queries without requiring controllers to set the total privacy loss budget. We demonstrate our contributions through an IRB-approved user study and experimental evaluations that show the PRI is useful for helping controllers choose $\epsilon$ and our algorithms are efficient. Overall, our work contributes to making DP easier to use for controllers by lowering adoption barriers.
Differential privacy, data controller, privacy loss budget
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data management systemsccs: Security and privacy Privacy protectionsccs: Security and privacy Usability in security and privacy
1. Introduction
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006) is a promising tool to analyze data about individuals, but it remains challenging to use in practice. In a typical DP deployment, a data analyst runs queries on a dataset containing sensitive information about individuals that is accessible to a data controller. We define a data controller as an agent with the capacity to (1) directly access and run analysts’ queries on individuals’ raw data (2) make decisions on what query output to release to the analysts since the output may reveal the identity of individuals. The controller is responsible for protecting individuals’ privacy while releasing useful output to the analysts. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: larger $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a major challenge exists which prevents broader DP deployment.
Challenge of choosing $\epsilon$. Controllers need to choose the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.  Dwork et al. (Dwork et al., 2019) showed that controllers have no clear consensus in choosing $\epsilon$ or approaching this problem in a principled manner.
Our Contributions
Designing PRI and the Find-$\epsilon$-from-PRI algorithm. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. We hypothesize knowing the effect of choosing $\epsilon$ on each individual’s privacy would help controllers make the choice. Therefore, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. The PRI is a relative indicator of ex-post per-instance privacy loss (Redberg and Wang, 2021), which is the realized privacy loss incurred to an individual upon releasing a specific output. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects each individual’s privacy. And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of PRIs consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release query output given a simple description of controllers’ privacy preferences.
Privately releasing $\epsilon$. The Find-$\epsilon$-from-PRI algorithm only releases the DP output and not the chosen $\epsilon$, since $\epsilon$ is chosen based on the PRIs, which are data-dependent, and releasing $\epsilon$ directly may undermine the DP guarantees. Thus we introduce a modification of the algorithm, called Find-and-release-$\epsilon$-from-PRI, which permits releasing both the query output and $\epsilon$ while preserving DP.
Bounding privacy loss across multiple queries. In many cases, analysts submit multiple queries to the controller to obtain different information about the dataset. Besides the challenge of choosing $\epsilon$ for each query, controllers also face the challenge of setting the global privacy budget which bounds the maximum privacy loss for a dataset. We propose a solution that does not require controllers to quantify the maximum acceptable privacy loss; instead, it dynamically bounds the privacy loss after answering (or denying) each query. We construct a privacy odometer (Rogers et al., 2016) to provide an end-to-end DP guarantee for our solution.
Evaluation. In our evaluation, we test the hypothesis that the new privacy risk indicator helps controllers choose $\epsilon$ by conducting an IRB-approved user study with 56 participants. We show that our algorithms are efficient and scalable, and we conduct ablation studies to analyze the effect of varying controllers’ privacy preferences.
Outline. The rest of the paper is organized as follows. We introduce the background of differential privacy and the different agents in standard DP deployment in Section 2. We present the new privacy risk indicator in Section 3 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 4. In Section 5, we introduce the modified algorithm to privately release $\epsilon$ and our solution to dynamically bound privacy loss across multiple queries. We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
2. Background
In this section, we first introduce the definition of $(\epsilon,\delta)$-differential privacy (Section 2.1), then explain the different agents in a standard DP deployment (Section 2.2).
2.1. Differential Privacy
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual. Formally, for an $\epsilon,\delta\geq 0$, a randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O)+\delta.$
When the failure parameter $\delta$ is 0, $\mathcal{M}$ satisfies $\epsilon$-differential privacy; otherwise $\delta$ is usually set to be a very small probability (e.g., $\frac{1}{n^{2}}$ where $n$ is the number of records). One way of achieving DP is by constructing a mechanism that adds random noise to the query output. The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$. Smaller $\epsilon$ leads to more noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff in DP (Near et al., 2021).
There are two important properties of differential privacy, post processing and sequential composition, which are building blocks for designing sophisticated mechanisms that satisfy DP.
Post processing (Dwork et al., 2014). DP is immune to post processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees. Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism. For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
Sequential composition (Dwork et al., 2014). Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy. Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
2.2. Agents in DP Deployment
Figure 1. Agents in standard DP deployment
Figure 1 shows a diagram of a standard DP deployment, i.e., the central DP model (Near et al., 2021) (Wagh et al., 2021); we define a dataflow as a function that passes data from one entity to another. Individuals send their raw data to a data controller. The data analyst submits queries to the controller in order to obtain information about the individuals. The goal of the controller is to release useful output to the analyst while protecting individuals’ privacy. To achieve this, the controller answers the analyst’s queries using DP mechanisms, and they control both how much protection they want to offer to individuals and how accurate the output should be by setting the $\epsilon$ parameter when applying DP to answer each query. In addition, the controller needs to set a global privacy budget to control the maximum acceptable privacy loss incurred on the dataset; answering each query consumes some budget, once the budget is reached, the dataset is retired and can no longer be queried.
One notable example of the central DP model is the application of DP by the US Census in 2020 (Abowd, 2018). In this case, the US Census Bureau is the data controller who has access to individuals’ (i.e., the US population) data. The data analysts’ queries are pre-determined by the Census Bureau; the census data is the output of various census queries (e.g., population counts in different geographical levels by age, ethnicity, race, and sex). The Census Bureau sets the privacy loss budget $\epsilon$ and allocates the budget to each query (Bureau, 2021b)  [^1] Technically, the Census Bureau allocates the budget using the privacy parameter $\rho$ in zero-concentrated differential privacy (Bun and Steinke, 2016), which can be converted to the $\epsilon$ in $(\epsilon,\delta)$-differential privacy. We omit the details for simplicity., then uses DP to release the query output. The Census Bureau made considerable effort to choose $\epsilon$ and allocate the budget, including testing different $\epsilon$ values on a demonstration census dataset and seeking extensive feedback from experts (Bureau, 2021a). However, when deploying DP in practice, choosing $\epsilon$ is an extremely challenging task for controllers without the assistance of DP experts (Dwork et al., 2019).
While there are other DP deployment models available, i.e., the local DP (Kasiviswanathan et al., 2011) or shuffler model (Erlingsson et al., 2019), where the agents interact differently (e.g., individuals no longer send their raw data to the controller), in this paper we concentrate on addressing the controller’s challenges of choosing $\epsilon$ in the central model.
3. Privacy Risk Indicator (PRI)
3.1. Why Choosing $\epsilon$ is Hard and PRI Overview
Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy. The parameter $\epsilon$ bounds the maximum privacy loss for any individual a priori before running the DP mechanism to produce the query output, but it is challenging for controllers to reason about the concrete privacy implications for individuals represented in the dataset when releasing a specific output ex-post, i.e., after applying the DP mechanism.
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output. The PRI is an extension to a more fine-grained characterization of the privacy loss in standard DP, thus it does not depart from or replace the standard DP theory. We construct PRI so that it is directly connected with the concept of ex-post per-instance privacy loss (Redberg and Wang, 2021), which we explain next in Section 3.2. This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$. One strategy includes an algorithm to find $\epsilon$ based on the controller’s privacy preferences expressed in terms of PRIs, which we will explain in Section 4.
3.2. Privacy Risk Indicator: Formal Definition
We represent a dataset as a set of $n$ records $x=(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, drawn from a data universe $\mathcal{X}$. WLOG, we assume each individual (labeled $1$ to $n$) contributes exactly one record to $x$.
We denote an $(\epsilon,\delta)$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon,\delta)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and $(\epsilon,\delta)$ as input, we define the privacy risk indicator of an individual $i\in[n]$ as follows:
Definition 0 (Privacy Risk Indicator).
The privacy risk indicator of an individual $i$ is defined as:
$PRI_{i}=\lVert r-q(x_{-i})\rVert_{v},$
where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$. The value of $v$ depends on the specific mechanism $\mathcal{M}$: if $\mathcal{M}$ is the Laplace Mechanism (Dwork et al., 2014), $v=1$ (i.e., the $\ell_{1}$ norm); if $\mathcal{M}$ is the Gaussian Mechanism (Dwork et al., 2014), $v=2$ (i.e., the $\ell_{2}$ norm).
We note that one limitation of our formulation of PRI is that it only applies to numeric queries, but not categorical queries that return an element from a set, since both $r$ and $q(x_{-i})$ must be in $\mathbb{R}^{k}$. One interesting direction is to explore a more generic formulation of PRI that adapts to both types of queries.
Connection between PRI and privacy loss. Next, we show the connection between PRI and ex-post per-instance privacy loss (Redberg and Wang, 2021), which is the privacy loss conditioned on a fixed output, an individual, and a dataset. Specifically, we show that when the Laplace Mechanism (Dwork et al., 2014) or the Gaussian Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of ex-post per-instance privacy loss.
Definition 0 (Ex-post per-instance privacy loss (Redberg and Wang, 2021)).
For a randomized mechanism $\mathcal{M}$ and a query $q\in\mathcal{Q}$, the ex-post per-instance privacy loss given a realized query output $r\sim\mathcal{Y}$, a fixed dataset $x$, and an individual’s record $x_{i}\in x$ is
$\epsilon(r,x,x_{-i})=\left|\ln\left(\frac{\Pr[\mathcal{M}(x,q)=r]}{\Pr[\mathcal{M}(x_{-i},q)=r]}\right)\right|$
Note the difference between $\epsilon$, which is used as input to a DP mechanism, and $\epsilon(\cdot)$, which is a function that denotes the privacy loss conditioned on some given input, i.e., ex-post per-instance privacy loss. Intuitively, using ex-post per-instance privacy loss answers the question: what is the realized privacy loss incurred to each individual in the dataset when $r$ is released? Note that the ex-post per-instance privacy loss is a more fine-grained characterization of the privacy loss in standard DP; if $\epsilon(\cdot)\leq\epsilon$ for all $r\subseteq Range(\mathcal{M}),x\in\mathcal{X}^{n}$, and $x$’s neighboring dataset $x^{\prime}$ where $\lVert x-x^{\prime}\rVert_{1}\leq 1$, then $\mathcal{M}$ satisfies $\epsilon$-DP (similar argument can be made for $(\epsilon,\delta)$-DP). We derive our notion of PRI from the ex-post per-instance privacy loss so that it describes the realized individual privacy loss when an output of a DP mechanism is released. The PRI works as a relative indicator of privacy loss for both Laplace Mechanism and Gaussian Mechanism, which we explain next.
Definition 0 (Laplace Mechanism (Dwork et al., 2014)).
Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon>0$, the Laplace Mechanism is defined as:
$\mathcal{M}_{L}(x,q,\epsilon)=q(x)+(Z_{1},\ldots,Z_{k}),$
where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{\Delta_{1}}{\epsilon}$  [^2] The global sensitivity of q in the $\ell_{1}$ norm is defined as $\Delta_{1}=\max\lVert q(x)-q(x^{\prime})\rVert_{1}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$.. The Laplace Mechanism satisfies $\epsilon$-DP.
When $r$ is released using the Laplace Mechanism, the ex-post per-instance privacy loss is computed as:
$\epsilon_{L}(r,x,x_{-i})$ $=\left|\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/\Delta_{1})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/\Delta_{1})}\right)\right|$
$=\left|\epsilon\frac{\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1}}{\Delta_{1}}\right|.$
Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
$\epsilon_{L}(r,x,x_{-i})>\epsilon_{L}(r,x,x_{-j})$
The proof is straightforward: $PRI_{i}>PRI_{j}$ means $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$ (the PRI is in $\ell_{1}$ norm since Laplace Mechanism is used), thus the inequality holds.
Definition 0 (Gaussian Mechanism (Dwork et al., 2014)).
Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon,\delta>0$, the Gaussian Mechanism is defined as:
$\mathcal{M}_{G}(x,q,\epsilon,\delta)=q(x)+(Z_{1},\ldots,Z_{k}),$
where $Z_{i}$ are i.i.d random variables drawn from the Gaussian distribution $\mathcal{N}(\sigma^{2})$ with center 00 and variance $\sigma^{2}=\frac{2\Delta_{2}^{2}\ln(1.25/\delta)}{\epsilon^{2}}$  [^3] The global sensitivity of q in the $\ell_{2}$ norm is defined as $\Delta_{2}=\max\lVert q(x)-q(x^{\prime})\rVert_{2}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$.. The Gaussian Mechanism satisfies $(\epsilon,\delta)$-DP.
When $r$ is released using the Gaussian Mechanism, the ex-post per-instance privacy loss is computed as:
$\epsilon_{G}(r,x,x_{-i})$ $=\left|\ln\left(\frac{\exp(-\lVert r-q(x)\rVert_{2}^{2}/2\sigma^{2})}{\exp(-\lVert r-q(x_{-i})\rVert_{2}^{2}/2\sigma^{2})}\right)\right|$
$=\left|\frac{\lVert r-q(x_{-i}\rVert_{2}^{2}-\lVert r-q(x)\rVert_{2}^{2}}{2\sigma^{2}}\right|.$
Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
$\epsilon_{G}(r,x,x_{-i})>\epsilon_{G}(r,x,x_{-j})$
The proof is similar to that of the Laplace Mechanism, except now the PRI is in $\ell_{2}$ norm. Hence $PRI_{i}>PRI_{j}$ means $\lVert r-q(x_{-i})\rVert_{2}$ $\lVert r-q(x_{-j})\rVert_{2}$.
We now make several remarks:
PRI gives controllers additional information to decide what to release. The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output. The PRI is private information available to controllers only. In this paper, we show how PRI helps controllers decide which output to release.
PRI is data-dependent and thus leaks information about individuals. Controllers already have access to individuals’ raw data, so when they gain access to the PRIs, they do not learn anything extra about the individuals. Similar to how controllers would not release the individuals’ raw data, they will not release the PRIs. PRIs must remain private to the controllers and never be shared with any other agents.
Controllers may give PRI different uses. Controllers may express their privacy preferences for individuals in the dataset in terms of their associated PRIs, and derive an $\epsilon$ that produces that set of PRIs. This means that eliciting controllers’ privacy preferences based on PRIs paves the way for choosing $\epsilon$. We exploit this insight in the next section to design an algorithm to find $\epsilon$. We also discuss other ways in which PRIs may be used in Section 4.3.
4. Deriving Epsilon based on the PRI
In this section, we describe the intuition of using PRIs to express controllers’ privacy preferences (Section 4.1), then introduce an algorithm to choose $\epsilon$ by using PRIs as a more intuitive proxy (Section 4.2). We discuss the algorithm’s complexity and extensions in Section 4.3.
4.1. Using PRI to Express Privacy Preferences
Traditionally, controllers need to choose $\epsilon$ to convey their privacy preferences. However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset. In contrast, the per-individual PRI indicates the realized privacy loss of an individual relative to other individuals in the dataset. The consequence is that the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output. Next, we convey the intuition on how the PRI is useful to controllers through a example.
Figure 2. Example of using PRI to find $\epsilon$. The query is to compute the average GPA of the three students. For each $\epsilon$ (including when no DP is used), we show a specific query output and the corresponding PRI of each student.
Example illustrating one use of the PRI. Figure 2 shows the resulting PRIs of three students after answering a query about their average GPA. If the query output were released without using DP (a), the analyst observing the output would learn much about those students’ grades, as indicated by the PRI (recall that the PRI is a proxy to privacy loss). If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism. This would lead to a noisy output and a PRI associated with each student (note that running the mechanism two times with the same $\epsilon$ does not guarantee the same PRIs are obtained). We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students. In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others (they have the same privacy loss). Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.275) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve. Lastly, using a smaller $\epsilon=0.1$ (d) also makes the PRIs roughly similar, but at the cost of producing a query output with much worse accuracy than (c). One sensible choice is to choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
The above is an illustrative example that does not work in practice, since finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
4.2. The Find-$\epsilon$-from-PRI Algorithm
When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the realized privacy loss of each individual compared to others if the query output is released. Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a profile of PRIs that represents their preferences. Rather than choosing $\epsilon$ and releasing the DP output, controllers first check the corresponding PRIs to understand how would releasing the output affect each individual’s privacy, and then they decide whether or not to release the output based on their privacy preferences on the individuals. We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI. In Section 4.1, we have shown one criterion for choosing $\epsilon$ that leads to exactly equal PRIs, which is typically infeasible. Instead, we propose an alternative criterion to let controllers decide whether a query output respects their privacy preferences. We explain the key idea and then the algorithm.
Key Idea. We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset. This is done by requesting exactly one parameter from the controllers, a percentage $p$ ($0\leq p\leq 100$), which is easier to interpret than $\epsilon$. We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$. Then we find an $\epsilon$ that produces a DP output with its corresponding PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$. In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI. For instance, setting $p=0$ means the PRIs should be exactly equal.
Unlike $\epsilon$, the parameter $p$ directly conveys how close the individuals’ realized privacy loss should be — the PRI is a relative indicator of ex-post privacy loss for each individual. Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$. The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$, because if controllers’ privacy preferences can be satisfied, larger $\epsilon$ is preferred as it leads to more accurate output.
Input : DP mechanism $\mathcal{M}$; dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
Output : Query output $r$
1
2/* Initialize set of $\epsilon$ values to test */
3 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
4
5/* Compute $q(x_{-i})$ for all individual $i$ using memoization */
6 $C\leftarrow map()$
7 for $x_{i}\in x$ do
8       if $x_{i}\notin C$ then
9             $C[x_{i}]\leftarrow q(x_{-i})$
10
11
12for $\epsilon\in\mathcal{E}$ do
13
14      /* Compute query output using DP with current $\epsilon$/
15       $r\leftarrow\mathcal{M}(x,q,\epsilon,\delta=\frac{1}{n^{2}})$
16
17      /* Compute corresponding PRI of each individual */
18       $PRIs\leftarrow list()$
19       for $x_{i}\in C$ do
20             if $\mathcal{M}$ is ’Laplace’ then
21                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
22            if $\mathcal{M}$ is ’Gaussian’ then
23                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{2}$
24            $PRIs.add(PRI_{i})$
25
26      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
27
28      /* Check whether PRIs satisfy controller’s privacy preference */
29       if $PRI_{min}/PRI_{max}\geq 1-\frac{p}{100}$ then
30             return $r$
31
return nil
Algorithm 1 Find-$\epsilon$-from-PRI
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output $r$. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ values is infinite, we set an upper bound and a range, e.g., 10 and 0.1, to ensure a finite set of options, e.g., 100 in this example. Then, for each $x$’s neighboring dataset $x_{-i}$, the algorithm computes $q(x_{-i})$ (line 1 - 1), which will be used to compute the PRI for each individual $i$. Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1). For each $\epsilon$, the algorithm computes the query output $r$ by applying the given DP Mechanism $\mathcal{M}$ with parameter $\epsilon$ and a fixed $\delta$ (line 1), then computes the corresponding PRI of each individual (line 1 - 1). Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1). If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output. Note that at this point $\epsilon$ is not released to the analysts, and only the controller knows which $\epsilon$ is chosen by the algorithm.
Figure 3. Example of $\epsilon$ chosen by the Find-$\epsilon$-from-PRI algorithm with different percentage $p$ over 50 runs
Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRIs across executions of the mechanism; simultaneously, similar PRIs may be produced by different $\epsilon$ values. What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$.
In addition, we show empirically in Section 6.3.1 that stricter privacy preferences (i.e., smaller $p$) lead to smaller $\epsilon$, as expected. For demonstration, we present one of our evaluation results (Q4 in Figure 10). Figure 3 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs. As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$. Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 0.1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
Figure 4. Dataflow of Find-$\epsilon$-from-PRI Algorithm
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 4 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm to find $\epsilon$ and release query output to the analyst. Individuals send their data to the controller; this dataflow remains unchanged from the traditional central DP model (Near et al., 2021). The analyst submits a query to the controller. The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to release a suitable query output, which is an output produced by a DP mechanism (Laplace Mechanism or Gaussian Mechanism in our work) with an $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The PRI associated with each individual is never released to the analyst. The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval. The value for $\epsilon$ that led to the query output is also not released since $\epsilon$ is chosen based on PRIs, which are data-dependent, so releasing it directly might jeopardize the DP guarantee. In Section 5, we explain a modification of the algorithm that permits releasing $\epsilon$ differential privately.
Claim: Algorithm 1 is $(\epsilon,\delta)$-DP, where $\epsilon$ is the value found by the algorithm that satisfies the PRI threshold, or 0 if no suitable $\epsilon$ is found.
Algorithm 1 only releases the query output $r$, which is computed using an $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when $\mathcal{M}$ is Laplace Mechanism and $\delta=\frac{1}{n^{2}}$ when $\mathcal{M}$ is Gaussian Mechanism). If no suitable $\epsilon$ is found, the algorithm does not release $r$. Therefore, the algorithm satisfies $(\epsilon,\delta)$-DP.
4.3. Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel. Thus, we implement parallelism and compute the query output for a chunk of records in each process, and we show in Section 6.2 that the algorithm is both efficient and scalable.
Expressing privacy preferences. By default, the algorithm considers the list of PRIs for all individuals in the dataset when finding $\epsilon$. A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of records. For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x^{\prime}$) and those younger ($x^{\prime\prime}$), and specify that the two groups need similar privacy protection. This preference can easily be incorporated into the algorithm by letting the controller change the two PRI values to compare instead of the global $PRI_{min}$ and $PRI_{min}$ for all individuals; e.g., they can compute $PRI_{med}^{\prime}$ and $PRI_{med}^{\prime\prime}$, the median PRI for individuals in $x^{\prime}$ and $x^{\prime}$ respectively, and choose an $\epsilon$ that makes $PRI_{med}^{\prime}/PRI_{med}^{\prime\prime}\geq 1-(p/100)$.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP. The PRI complements these existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
5. Releasing $\epsilon$ and Bounding Privacy Loss
In Section 5.1, we propose a modification of Find-$\epsilon$-from-PRI that finds $\epsilon$ and releases both the query output and $\epsilon$ simultaneously while preserving DP. In Section 5.2, we discuss how to bound the total privacy loss in the presence of multiple queries and we provide the end-to-end DP guarantee in Section 5.3.
5.1. Privately Releasing $\epsilon$
Since Algorithm 1 finds $\epsilon$ in a data-dependent manner, releasing the chosen $\epsilon$ directly may reveal additional information about the dataset and violate the DP guarantees. Therefore, a modification of the algorithm is needed to ensure DP is preserved if the controller wants to release both the query output $r$ and $\epsilon$ to the analysts.
Key Idea. Our key idea is to leverage the Sparse Vector Technique (SVT) (Dwork et al., 2009) to test multiple $\epsilon$ values (from largest to smallest) and release the first $\epsilon$ that satisfies the PRI threshold specified by the controller. Algorithm 1 tests the set of $\epsilon$ values and chooses the largest $\epsilon$ that passes the PRI threshold; by incorporating SVT, the modified algorithm chooses approximately the largest $\epsilon$ that passes the threshold so releasing it preserves DP.
Input : Dataset $x$; a stream of SVT queries $\mathcal{Q}_{svt}=q_{1},q_{2}\cdots$ with global sensitivity at most $\Delta_{svt}$; threshold $\tau$; privacy budget $\epsilon_{svt}^{\prime}$ and $\epsilon_{svt}^{\prime\prime}$
Output : A stream of answers $a_{1},a_{2}\cdots$ where each $a_{i}\in\{\top,\bot\}$
1
[…]
3for $q_{i}\in\mathcal{Q}_{svt}$ do
4
5      if $q_{i}(x)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq\tau+\rho$ then
6             Output $a_{i}=\top$
7             return
8      else
9             Output $a_{i}=\bot$
10
Algorithm 2 SVT algorithm proposed by Lyu et al. (Lyu et al., 2016)
The SVT algorithm. Algorithm 2 shows the standard SVT algorithm proposed by  Lyu et al. (Lyu et al., 2016)  [^4] The algorithm from  (Lyu et al., 2016) has an additional parameter $c$ (the maximum number of queries the algorithm can answer $\top$). We set $c=1$, since we only need one positive answer for our modified algorithm., which we will adopt and incorporate into our modified algorithm to release $\epsilon$ privately. SVT takes a dataset, a stream of queries (each with a global sensitivity of at most $\Delta_{svt}$), a given threshold, and a privacy budget $\epsilon_{svt}$ that is divided into $\epsilon_{svt}^{\prime}$ and $\epsilon_{svt}^{\prime\prime}$ (i.e., $\epsilon_{svt}=\epsilon_{svt}^{\prime}+\epsilon_{svt}^{\prime\prime}$), then outputs a stream of answers indicating whether each query output is above or below the threshold. The algorithm works by first initializing a Laplacian noise ($\rho$) scaled to $\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}$ (line 2). For each SVT query, it generates another term of Laplacian noise scaled to $\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}$ and adds the noise to the query output. Then the algorithm compares the noisy query output with the noisy threshold (line 2); if it is above or equal to the threshold, the algorithm outputs $\top$ indicating the query output approximately passes the threshold, otherwise, it outputs $\bot$. Note that the algorithm answers the SVT queries in a streaming fashion. Overall, the algorithm satisfies $\epsilon_{svt}$-DP.
Input : DP mechanism $\mathcal{M}$; dataset $x=(x_{1},\cdots,x_{n})$; query $q$; variance threshold $\tau$; SVT budget $\epsilon_{svt}$
Output : Query output $r$, chosen $\epsilon$
1
2/* Assign privacy budget for running SVT */
3 $\epsilon_{svt}^{\prime},\epsilon_{svt}^{\prime\prime}\leftarrow\frac{\epsilon_{svt}}{1+2^{2/3}},\frac{2^{2/3}\epsilon_{svt}}{1+2^{2/3}}$
4
5/* Global sensitivity of variance */
6 $\Delta\leftarrow\frac{1}{n}$
7
8/* Generate noise to add to PRI threshold */
9 $\rho=Lap\left(\frac{\Delta}{\epsilon_{svt}^{\prime}}\right)$
10
11/* Initialize set of $\epsilon$ values to test */
12 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
13
14/* Compute $q(x_{-i})$ for each individual $i$ using memoization */
15 $C\leftarrow map()$
16 for $x_{i}\in x$ do
17       if $x_{i}\notin C$ then
18             $C[x_{i}]\leftarrow q(x_{-i})$
19
20
21for $\epsilon\in\mathcal{E}$ do
22
23      /* Compute query output using DP with current $\epsilon$/
24       $r\leftarrow\mathcal{M}(x,q,\epsilon,\delta=\frac{1}{n^{2}})$
25
26      /* Compute corresponding PRI of each individual */
27       $PRIs\leftarrow list()$
28       for $x_{i}\in C$ do
29             if $\mathcal{M}$ is ’Laplace’ then
30                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
31            if $\mathcal{M}$ is ’Gaussian’ then
32                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{2}$
33            $PRIs.add(PRI_{i})$
34
35      /* Normalize PRIs and compute variance */
36       $PRI_{max}\leftarrow\max(PRIs)$
37       for $i\in range(PRIs.length)$ do
38             $PRIs[i]\leftarrow\frac{PRIs[i]}{PRI_{max}}$
39      $Var\leftarrow$ compute-variance($PRIs$)
40
41      /* Using SVT to release the output $r$ and $\epsilon$ that approximately passes variance threshold */
42
43      if $-Var+Lap\left(\frac{2\Delta}{\epsilon_{svt}^{\prime\prime}}\right)\geq-\tau+\rho$ then
44             return $r,\epsilon$
45
46
return nil
Algorithm 3 Find-and-release-$\epsilon$-from-PRI
What is a good SVT query for choosing $\epsilon$? To incorporate SVT in our algorithm, each SVT query corresponds to a specific $\epsilon$ the algorithm is testing, and each query should output a value based on PRIs computed from the $\epsilon$; naturally, this value could be the existing PRI criterion we already have in Algorithm 1, which is $\frac{PRI_{min}}{PRI_{max}}$. Then the algorithm applies SVT by comparing the PRI criterion to the threshold (i.e., the percentage $p$ specified by controllers) and releasing $\epsilon$ of which the corresponding PRI criterion approximately passes the threshold. However, the current PRI criterion has one caveat that makes it difficult to incorporate with SVT: the SVT query, which is a function that outputs $\frac{PRI_{min}}{PRI_{max}}$, has a sensitivity $\Delta_{svt}=1$. This means the algorithm will be adding a Laplacian noise scaled to $\frac{2}{\epsilon_{svt}^{\prime\prime}}$ to $\frac{PRI_{min}}{PRI_{max}}$; clearly, when $\epsilon_{svt}^{\prime\prime}$ is small, the magnitude of added noise will be too large and make the PRI criterion useless since it will be too noisy to choose a reasonable $\epsilon$.
Using variance as the new PRI criterion. To address this caveat, we opt for a different PRI criterion when using SVT — by using the variance of normalized PRIs. Each SVT query will output the variance of PRIs for the corresponding $\epsilon$, and the variance is compared to a threshold specified by the controller; if the variance is approximately smaller than the threshold, then a suitable $\epsilon$ is found. The normalized PRIs will have values between 0 and 1 — this still preserves the PRI’s meaning as a relative indicator of ex-post per-individual privacy loss. The variance of the PRIs will have a sensitivity of $\frac{1}{n}$, where n is the number of records in the dataset; this means when $n$ is relatively large, the sensitivity of the variance (hence the Laplacian noise added to the variance) will be small! In addition, the variance of PRIs shares the same intuitiveness as $\frac{PRI_{min}}{PRI_{max}}$: if the controllers want the individuals to have similar realized privacy loss (hence more protected), then they should set a lower variance threshold, and vice versa.
The Find-and-release-$\epsilon$-from-PRI algorithm. We present our modified Find-and-release-$\epsilon$-from-PRI algorithm in Algorithm 3. The main modification lies in line 3 -  3 and line 3 -  3. In line 3, the algorithm allocate the SVT budget $\epsilon_{svt}$ to have $\epsilon_{svt}^{\prime}:\epsilon_{svt}^{\prime\prime}=1:2^{\frac{2}{3}}$; we choose this allocation strategy from  (Lyu et al., 2016) as it leads to better accuracy of SVT. Then the algorithm iterates over each $\epsilon$ and computes its corresponding PRIs the same way as in Algorithm 1 (line 3 -  3). It then normalizes each PRI and computes the variance (line 3), and lastly, it compares the noisy variance with the noisy threshold (controllers want $Var\leq\tau$, which is equivalent to $-Var\geq-\tau$, and returns the current output $r$ and $\epsilon$ if it passes the threshold.
The main advantage of using SVT is that it always consumes a fixed budget $\epsilon_{svt}$ regardless of the number of SVT queries. Controllers understand the effect of using a specific $\epsilon_{svt}$ and running SVT to release $\epsilon$, since they know exactly how much noise is added to both the variance and the variance threshold $\tau$ they have specified. They can decide whether they are comfortable releasing an output that does not exactly conform to their privacy preferences; if not, they may use a different $\epsilon_{svt}$ or $\tau$ and rerun the algorithm.
Claim: Algorithm 3 is $(\epsilon+\epsilon_{svt},\delta)$-DP, where $\epsilon$ is the value released using SVT, or 0 if no suitable $\epsilon$ is found.
Algorithm 3 can be divided into two steps. In the first step, it uses a $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when the mechanism is Laplace) to compute the query output $r$. In the second step, it uses SVT, which is $\epsilon_{svt}$-DP, to output whether the SVT query for the given $\epsilon,r$ is above the threshold, and releases $\epsilon,r$ if it is above the threshold. Therefore, by sequential composition (Dwork et al., 2014), Algorithm 3 satisfies $(\epsilon+\epsilon_{svt},\delta)$-DP.
5.2. Bounding Privacy Loss Across Queries
The preceding discussion is valid for one query. In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery (Dwork et al., 2014). In today’s deployment of DP, controllers need to set a total privacy budget on the dataset a priori and stop accepting queries when the cumulative $\epsilon$ exceeds the budget (Dwork et al., 2019). The need for setting a total budget poses a critical challenge for controllers:
Controllers do not know the privacy implications of total privacy budget. Similar to choosing $\epsilon$ for a single query, when deciding which total budget to set, controllers face the same challenge of not knowing the actual privacy implications on individuals after releasing multiple query outputs. Oftentimes, they resort to selecting the total budget arbitrarily or through trial-and-error (Dwork et al., 2019).
The solution. Instead of quantifying the maximum acceptable privacy loss, which is difficult, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ or $\epsilon+\epsilon_{svt}$ of all previously answered queries for Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI respectively; we call it $\epsilon_{c}$. For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). The algorithm rejects the query if it fails to find a suitable $\epsilon$ larger than $\epsilon_{c}$, meaning the controller’s privacy preferences cannot be met for any possible $\epsilon$.
Our solution creates a dynamic bound of total privacy loss. As $\epsilon_{c}$ keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since larger $\epsilon$ may not produce PRIs that satisfy controllers’ privacy preferences. At each point the algorithm fails to find a suitable $\epsilon$ for a query, the total privacy loss is bounded. The bound is refreshed if a suitable $\epsilon$ is found for a different query. A potential concern is whether there is additional privacy leakage using this dynamic bound. In Section 5.3, we show the bound can be formalized as a privacy odometer in fully-adaptive composition (Rogers et al., 2016) and there is no leakage.
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ or $\tau$ in each algorithm based on analysts’ preferences, e.g., they can set a higher $p$ or $\tau$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ or $\tau$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
5.3. End-to-End DP Guarantee
Implications of choosing $\epsilon$ adaptively. If the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ or $\epsilon+\epsilon_{svt}$ for Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI respectively. However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$. Then the total privacy loss is not bounded by any fixed constant since the algorithm does not require setting a total privacy budget. This brings the question of whether releasing multiple query outputs computed using adaptively chosen $\epsilon$ causes additional privacy leakage.
Privacy odometer in fully-adaptive composition. To answer this question, we construct a privacy odometer (Rogers et al., 2016), which provides a probabilistic running upper bound on the realized privacy loss after answering each query. The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
In short, there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition. Therefore, the privacy odometer for Find-$\epsilon$-from-PRI under sequential composition (Dwork et al., 2014) is defined as:
Definition 0 (Privacy Odometer).
[…]
The privacy odometer for Find-and-release-$\epsilon$-from-PRI can be defined using the same principle, which we omit. Essentially, the privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}=\sum_{i=1}^{i=k}\epsilon_{i}$ after the algorithm answers the $k$-th query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $(\epsilon_{c},\delta_{g})$-DP (we can set $\delta_{g}=\sum_{i=1}^{i=k}\delta$).
It is also worth noting that under certain conditions (Whitehouse et al., 2023), the privacy odometer achieves the same asymptotic bound as advanced composition (Dwork et al., 2010). For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition, and we leave further improvements for future work.
6. Evaluation
We answer the following research questions:
  • RQ1: Does presenting the PRI help data controllers choose $\epsilon$?
  • RQ2: Are our proposed algorithms, Find-$\epsilon$-from-PRI and
    Find-and-release-$\epsilon$-from-PRI efficient and scalable?
  • Microbenchmarks: We study the effect of varying parameter $p$ in Find-$\epsilon$-from-PRI, and parameter $\tau$ in Find-and-release-$\epsilon$-
    from-PRI
    , and we compare the chosen $\epsilon$ values by both algorithms.
6.1. RQ1: Does PRI Help Controllers Choose $\epsilon$?
To determine whether PRI helps the data controllers choose $\epsilon$, we conducted an IRB-approved user study to analyze how PRI affects the controller’s decision-making regarding choosing $\epsilon$. To this end, we conducted a between-subjects user study where the participants are divided into control and treatment groups. To prevent learning effects, only the treatment group sees the PRI. We chose to conduct the study in the form of an online survey because our tasks do not require in-person guidance and it is more convenient for participants to complete the study. In addition, it allows us to collect data from a more representative sample.
6.1.1. Recruitment and data collection
We recruited 100 participants using Prolific (pro, [n. d.]). To complete the study, each participant must sign a consent form, which describes the purpose and content of the study, compensation, and data confidentiality. We do not collect identifiable information, and the survey responses are anonymized, meaning the participant’s IP address, location data, and contact information are not being recorded.
Prescreening criteria. When recruiting on Prolific, we prescreen participants based on three criteria: i) they are fluent in English, to avoid confusion due to misunderstanding language; ii) they have completed an undergraduate degree or above; iii) they study Computer Science, Computing (IT) or Mathematics; this is to ensure that the participants have a sufficient quantitative background to complete the study.
Tutorial and Screening Quiz. Since the main goal of the study is to ask the participant to choose $\epsilon$ when using DP to release query output, the participant needs to have basic knowledge about DP and understand the implications of the $\epsilon$ they chose. To ensure each participant knows enough about DP to complete the study, we ask each participant to read a short tutorial about DP at the beginning. This tutorial explains the intuition of DP and the effect of the parameter $\epsilon$ at a high level. Mainly, it explains that DP is generally achieved by adding random noise to the query output, and larger $\epsilon$ means less noise will be added and vice versa.
After reading the tutorial, the participant needs to complete a screening quiz that tests their knowledge about the tutorial they just read. Only participants who pass the quiz proceed to conduct the study. Of the 100 participants we recruited, 56 passed the quiz and their responses were used for the study analysis. This quiz ensures that participants have understood the principles of DP, thus strengthening the validity of the user study.
6.1.2. Study procedure
First, the participants are randomly divided into control or treatment groups; 29 participants are in the control group, and 27 are in the treatment group. The survey presents a sample from the UCI Adult dataset (uci, [n. d.]) and explains a DP mechanism is used to answer queries on the dataset by adding noise controlled by the parameter $\epsilon$. Then, it presents two queries in random order:
  • The number of people with income greater than 50K and are less than 40 years old.
  • The number of people who have a Master’s degree or above and have income $\leq$ 50k, grouped by each race.
For each query, the survey asks the participant to choose an $\epsilon$ value from a slider, with values from 0.5 to 5.0 in an interval of 0.5. Since we want all participants to have a clear task goal in mind, the survey includes a description of the task: choose $\epsilon$ to protect individuals’ privacy while yielding as accurate answers as possible.
For the control group, the participant is presented with a table showing the range of possible output for each $\epsilon$ to indicate output accuracy and the raw (non-DP) output. The table includes information that any controller would have access to in a deployment today when choosing $\epsilon$.
For the treatment group, the participant is presented with the same table, but with an additional column showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI); we also provide a brief explanation of the PRI. Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include it to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
6.1.3. Results
Figure 5. $\epsilon$ chosen by each participant
If the intervention (showing the PRIs) is effective, we expect this would manifest as more consistent answers among participants, since all participants were given the same training and the same task. Therefore, we compare the distribution of $\epsilon$ chosen by participants in the control and treatment groups.
Figure 5 shows a swarm plot where each point represents the $\epsilon$ value chosen by each participant in the control or treatment group. For both queries, 19 out of 27 participants in the treatment group chose either 2 or 1.5, showing a more consistent answer than the control group. This is because the PRI range when $\epsilon$ is 2 or 1.5 is considerably smaller than that of larger $\epsilon$, and smaller $\epsilon$ has lower accuracy. This shows that most participants could connect the PRIs and their task goal, which is to protect individuals and release useful output, and they were able to use the PRIs to choose the appropriate $\epsilon$ that satisfies the goal. On the contrary, the $\epsilon$ values chosen by participants in the control group are more spread out and have greater variance, suggesting participants shared no general consensus as to what is a good $\epsilon$ to choose.
We ran Mann–Whitney test  (McKnight and Najab, 2010) on the $\epsilon$ values chosen from the control and treatment group, with the alternative hypothesis being the two samples come from different distributions. For both queries, we obtained $p<0.005$. Thus we reject the null hypothesis and conclude that showing the PRIs helped participants choose $\epsilon$.
Last, we asked participants in the treatment group if they used the PRIs to complete the task: all 27 participants responded they did use that information, confirming they were indeed “treated”.
Figure 6. $\epsilon$ chosen by participants who do not know DP (left) and those who know DP (right)
Does knowing DP matter? Another interesting question is whether the participant’s prior familiarity with DP affects how they choose $\epsilon$. At the start of the survey before showing the tutorial, we asked each participant whether they know about DP. 27 participants (out of a total of 56) answered they do not know anything about DP and 29 participants answered they know about DP.
Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity. We find similar patterns. Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance. This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they knew about DP previously.
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$ instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and we asked them to complete the same survey questions as in the treatment group except that the survey presents histograms of PRIs instead of range plots.
Figure 7. $\epsilon$ chosen by participants who were presented with histograms or range plots, or in the control group
Figure 7 shows $\epsilon$ chosen by participants who were presented with histograms or range plots, and we include the $\epsilon$ chosen by participants in the previous control group as reference. Participants who saw histograms did not perform better than participants who saw range plots. Especially for Q2, eight participants chose larger $\epsilon$ (3.5, 4, and 4.5). Since the responses are anonymous, we cannot directly ask participants about their decision-making after the survey ends. We deduce one possible explanation might be that for the large $\epsilon$, the majority of PRIs are already small and similar, and only a small portion of PRIs have large values. The participants might think that as long as the privacy of the majority of people is being protected, the task goal is achieved; hence, they chose larger $\epsilon$, so the results are more accurate. However, if the participants only saw the range of PRI, they would be more inclined to choose smaller $\epsilon$ with a smaller range of PRI. Another possible explanation could be that some participants had difficulties interpreting histograms, whereas the range plots are generally much easier to understand.
The discrepancy of results by tweaking how we present the PRIs demonstrates how we convey information to the users affects their ability and reasoning for choosing $\epsilon$. Exploring other alternative solutions for presenting and helping users interpret PRI remains an interesting problem.
6.1.4. Threats to validity
External validity: We recruited participants from diverse backgrounds worldwide via Prolific, with few prescreen criteria so that the participants could complete the study without difficulty. Internal validity: To ensure participants have the minimum knowledge of DP needed to complete the study, each participant must read a tutorial and pass a quiz. Participants can only take the quiz once to minimize the chance of them abusing it if they have multiple chances. To avoid learning effects, we used a between-subject study to isolate the condition we intended to test. We also randomized the two questions for choosing $\epsilon$.
6.2. RQ2: Are Both Algorithms Practical?
Since both Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI algorithms can run multiple times if there are multiple queries, the algorithms need to be efficient so running it multiple times will still be practical. Thus, we evaluate both algorithms’ runtime for different queries and dataset sizes.
Datasets. We controlled the experiment by generating datasets with 1000, 10000, 100000, and 1 million records based on the Adult dataset (uci, [n. d.]). We sampled records for datasets of sizes (number of records) 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all attributes in the dataset.
Queries. We ran both algorithms for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.
  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num==13 AND age==25
  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==
    ’Asian-Pac-Islander’ AND 30<=age<=40 GROUP BY marital_status
  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=
    ’United-States’ AND sex==’Female’
  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass in (’Federal-gov’, ’Local-gov’, ’State-gov’)
  • Q5: SELECT SUM(capital_gain) FROM adult
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total). We set $p=5$ for Find-$\epsilon$-from-
PRI
, and $\epsilon_{svt}=1,\tau=10^{-6}$ for Find-and-release-$\epsilon$-from-PRI. We used eight parallel processes for computing the PRIs.
Configurations. We ran the experiments on a Chameleon Cloud (cha, [n. d.]) instance with 160 CPUs (Intel Xeon Platinum 8380) and 256GB memory. We used Ubuntu 22.04 and Python 3.10 for all experiments.
Figure 8. Average runtime (in log scale) to find $\epsilon$ over 10 runs by Find-$\epsilon$-from-PRI (a) and Find-and-release-$\epsilon$-from-PRI (b)
Results. Figure 8 shows two line plots where the x-axis is the data size and the y-axis is the runtime, both in log scale; the left plot corresponds to the run time of Find-$\epsilon$-from-PRI, and the right plot corresponds to the run time of Find-and-release-$\epsilon$-from-PRI. Each line in the plot represents a query, with the error band showing the standard deviation.
As shown in both plots, the runtime of both algorithms increases linearly with the data size. Most queries took less than a few minutes to finish, even for 1 million records. The runtime of Find-and-release-$\epsilon$-from-PRI is slightly longer than that of Find-$\epsilon$-from-PRI, especially when the dataset size is small. This is because applying SVT in Find-$\epsilon$-from-PRI takes an extra overhead since it needs to normalize the PRIs, compute the variance, and generate two Laplacian noise. However, when the dataset gets larger, the overhead becomes negligible and the runtime is dominated by the time to compute PRIs.
6.3. Microbenchmarks
Figure 9. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$. The mechanism used is Laplace.
Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$. The mechanism used is Gaussian.
6.3.1. Effect of parameter $p$ in Find-$\epsilon$-from-PRI
The Find-$\epsilon$-from
-PRI
algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$. Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer. We test the effect of varying $p$ in the algorithm; specifically, when $p=5,25,50,75,95$.
Setup. We use the Adult dataset with 100000 rows that we generated in Section 6.2. We ran each of the five queries from Section 6.2 50 times, and recorded the $\epsilon$ chosen by the algorithm in each run.
Results. Figure 9 and Figure 10 show the boxplots of $\epsilon$ found by the algorithm with different $p$ for each query; Figure 9 shows the results when the DP mechanism $\mathcal{M}$ used in the algorithm (for computing the query output $r$ and the PRIs) is Laplace Mechanism, and Figure 10 shows the results for the Gaussian mechanism. As shown in both figures, setting larger $p$ tends to find larger $\epsilon$. Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 when the mechanism is Laplace, and less than 1 when the mechanism is Gaussian. In fact, for most of the queries and $p$, using Gaussian caused the algorithm to choose larger $\epsilon$ compared to using Laplace, because the Gaussian Mechanism tends to produce noisier output than the Laplace Mechanism  [^5] This holds when the query’s $\ell_{2}$ sensitivity (used by Gaussian Mechanism) is similar to its $\ell_{1}$ sensitivity (used by Laplace Mechanism); otherwise, if the query’s $\ell_{2}$ sensitivity is much smaller, Gaussian Mechanism adds less noise compared to Laplace Mechanism. using the same $\epsilon$. In addition, both algorithms find larger $\epsilon$ when $p>50$ for queries with larger sensitivity (Q2, Q4, Q5), and they tend to find smaller $\epsilon$ for queries with sensitivity of 1 (Q1, Q3).
Overall, the results empirically demonstrate that both algorithms are useful for finding a suitable $\epsilon$ that is consistent with the controller’s privacy preferences.
Figure 11. $\epsilon$ found by Find-and-release-$\epsilon$-from-PRI algorithm over 50 runs, with $\tau=10^{-3},10^{-6},10^{-9},10^{-12}$.
6.3.2. Effect of parameter $\tau$ in Find-and-release-$\epsilon$-from-PRI
The Find-and-release-$\epsilon$-from-PRI algorithm lets controllers specify the PRI variance threshold $\tau$ (and the SVT budget $\epsilon_{svt}$), and it uses SVT to choose the first $\epsilon$ of which the variance of its corresponding PRIs is approximately smaller than the threshold. Smaller $\tau$ means stricter privacy preferences. We test the effect of varying $\tau$ in the algorithm; specifically, when $\tau=10^{-3},10^{-6},10^{-9},10^{-12}$.
Setup. We use the same setup as in Section 6.3.1. We set $\epsilon_{svt}=1$ and we use Laplace Mechanism as the DP mechanism $\mathcal{M}$.
Results. Figure 11 shows the boxplots of $\epsilon$ found by the algorithm with different $\tau$ for each query. The $\epsilon$ values selected are virtually similar and small when $\tau=10^{-6},10^{-9},10^{-12}$. On the contrary, the algorithm mostly selects the largest $\epsilon$ (10) it is considering when $\tau=10^{-3}$ for queries Q1, Q2, Q4, and Q5, and it selects a relatively small $\epsilon$ that is less 1 for Q3. This suggests controllers may need to observe the PRIs to decide a reasonable variance threshold to release $\epsilon$ and the query output $r$. As we noted before in Section 5.1, controllers may also rerun the algorithm multiple times with different parameters $\tau$ and $\epsilon_{svt}$ to observe how much noise is added and decide what to release that conforms to their privacy preferences.
7. Related Work
Privacy loss accounting. Various DP definitions have been proposed to account the privacy loss in a data-dependent manner or in different granularity. Our paper defines PRI as a relative indicator of the ex-post per-instance privacy loss (Redberg and Wang, 2021), which is built upon previous works that separately proposed ex-post privacy loss (Ligett et al., 2017) and per-instance privacy loss (Wang, 2019); ex-post privacy loss is the privacy loss conditioned on the realized output of a DP mechanism, and per-instance privacy loss is the privacy loss conditioned on a fixed dataset and a specific record in it. In addition,  Feldman and Zrnic (Feldman and Zrnic, 2021) proposed a privacy filter under Renyi differential privacy, which tracks the personalized privacy loss for each individual and drops an individual’s record when their personalized privacy loss exceeds a fixed budget (for all individuals). Seeman et al. (Seeman et al., 2024) proposed a formulation of DP called Per Record Differential Privacy; in this formulation, only the policy function on each individual’s privacy loss is published but not the actual privacy loss value, and individuals can compute their own privacy loss using the function. Unlike most works which define and use a different DP formation than standard DP, our work does not try to replace standard DP — we simply use the notion of ex-post per-instance privacy loss as a tool to connect our PRI definition to privacy loss in a more fine-grained granularity, and both our algorithms, Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI, satisfy $(\epsilon,\delta)$-DP.
Choosing privacy loss budget. Different approaches (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Hay et al., 2016; Ligett et al., 2017) aim to help users choose $\epsilon$ when using DP. One line of work (Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021) implements interfaces to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. (Nanayakkara et al., 2022) visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton (Lee and Clifton, 2011), which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ligett et al. (Ligett et al., 2017) empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting. Ge et al. (Ge et al., 2019) translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. (Hsu et al., 2014) proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts. Cummings et al. (Cummings et al., 2021) and Xiong et al. (Xiong et al., 2020) investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
Customizing Privacy. Some works study methods to allow users to specify privacy requirements at the individual level. Jorgensen et al. (Jorgensen et al., 2015) proposes Personalized Differential Privacy in which each individual specifies their own privacy preferences instead of using a global privacy loss budget. Since choosing a numerical privacy preference value is challenging for typical users, the authors recommended associating suitable privacy preference values to different groups (e.g., low/medium/high privacy). He et al. (He et al., 2014) proposes Blowfish Privacy in which in addition to specifying $\epsilon$, individuals specify their privacy policies consisting of secrets, i.e., information that must be kept secret about the individuals, and constraints, i.e., information that may be known publicly and exploited by adversaries. Both works attempt to define the granularity of privacy protection on the individual level but put the burden on the users to specify their privacy preferences. Instead, PRI is interpretable and controllers can specify their privacy preferences by observing the PRI of each individual directly.
8. Conclusions
In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain more understanding of the effects of applying differential privacy to the output of an analyst’s query. This translates into less burden in choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI. In addition, we propose an alternative algorithm, called Find-and-release
-$\epsilon$-from-PRI
, which finds and releases $\epsilon$ privately. We present solutions to bound privacy loss when answering multiple queries without requiring controllers to set a global privacy budget. Overall, our work reduces the burden for controllers to use DP in practice.